### v1

# Pinned distances of planar sets with low dimension

Abstract.
In this paper, we give improved bounds on the Hausdorff dimension of pinned distance sets of planar sets with dimension strictly less than one. As the planar set becomes more regular (i.e., the Hausdorff and packing dimension become closer), our lower bound on the Hausdorff dimension of the pinned distance set improves. Additionally, we prove the existence of small universal sets for pinned distances. In particular, we show that, if a Borel set $X\subseteq\mathbb{R}^{2}$ is weakly regular ($\dim_{H}(X)=\dim_{P}(X)$), and $\dim_{H}(X)>1$, then
$\sup\limits_{x\in X}\dim_{H}(\Delta_{x}Y)=\min\{\dim_{H}(Y),1\}$,
for every Borel set $Y\subseteq\mathbb{R}^{2}$. Furthermore, if $X$ is also compact and Ahlfors-David regular, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in X$ such that
$\dim_{H}(\Delta_{x}Y)=\min\{\dim_{H}(Y),1\}$.
The first author was supported in part by NSF DMS-2037851 and NSF DMS-2246906.
1. Introduction
Let $E\subseteq\mathbb{R}^{n}$, and define the distance set of $E$ to be the set
$\Delta(E):=\{|x-y|\mid x,y\in E\}$.
A fundamental problem in geometric measure theory is to relate the (Hausdorff) dimension of $E$ and the dimension of its distance set $\Delta(E)$. Falconer conjectured [7] that, if $E$ is analytic and $\dim_{H}(E)>n/2$, then $\Delta(E)$ has positive Lebesgue measure (and so $\dim_{H}(\Delta(E))=1$). This is still unproven in all dimensions $n\geq 2$.
A generalization of this is the pinned distance problem. If $E\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, the pinned distance set of $E$ with respect to $x$ is the set
$\Delta_{x}(E):=\{|x-y|\mid y\in E\}$.
The dimension version of Falconer’s conjecture for pinned distance sets is the following.
Conjecture 1.
Let $E\subseteq\mathbb{R}^{n}$ be an analytic set with $\dim_{H}(E)>n/2$. Then
$\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))=1$.
Substantial progress has been made on this question in the past few years [6, 9, 3, 4, 5].
A full resolution of Conjecture 1 problem seems to be beyond current techniques, even in the $n=2$ case. However, recent work has proven strong lower bounds on the dimension of pinned distance sets in the plane [23, 24, 11, 27]. If $E\subseteq\mathbb{R}^{2}$ is analytic, and $d:=\dim_{H}(E)>1$, the current records are
[…]
Unfortunately, the techniques in both [12] and [8] rely crucially on the assumption that $\dim_{H}(E)>1$.
In this paper, we prove lower bounds on the pinned distance sets of planar sets of small dimension. There is significantly less known in this setting. A fundamental issue had been the lack of progress on radial projections for low-dimensional sets. Nevertheless, Shmerkin and Wang gave the first strong progress on this problem [^1] A key ingredient of Shmerkin and Wang’s proof was a result on radial projections of low dimensional sets. This component was then improved in [22].. They proved [25] that, for any analytic set $E\subseteq\mathbb{R}^{2}$ with $d:=\dim_{H}(E)\leq 1$,
(1) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{d}{2}+\frac{d^{2}}{2(2+\sqrt{d^{2}+4})}.$
Moreover, they proved that if $E$ is weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$, then $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)=d$. In [5], Du, Ou, Ren and Zhang proved that
(2) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{5}{3}d-1,$
which improves on (1) when $d$ is sufficiently close to one.
In this paper, we establish stronger bounds on the dimension of pinned distance sets, for certain ranges of $d$. Our main theorem is the following.
Theorem 2.
Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set, $d:=\dim_{H}(E)\leq 1$ and $D:=\dim_{P}(E)$. Then,
$\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{\alpha D-d(D+\alpha-d)}{(d+1)(\alpha D-d^{2})-d^{2}(\alpha+D-2d)}\right)$,
where $\alpha=\min\{1+d,D\}$.
Note that the bound improves as $E$ becomes more regular, i.e., as the packing dimension of $E$ approaches its Hausdorff dimension. An important special case (when $D=2$) is captured by the following corollary.
Corollary 3.
Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set and $d:=\dim_{H}(E)\leq 1$. Then,
$\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{2-d}{2(1+2d-d^{2})}\right)$.
The proof of this theorem uses effective methods, and follows the general framework of the authors previous paper [8]. The assumption that $\dim_{H}(E)<1$ necessitates important changes to previous work.
One of the most crucial differences is the use of the recent breakthrough by Orponen, Shmerkin and Wang [22] on radial projections for low-dimensional sets. In the language of effective dimension, this allows us to find $x,y\in E$ whose direction $e:=\frac{y-x}{|x-y|}$ is sufficiently complex. Additionally, a number of the algorithmic tools of [8] need to be adapted to this setting.
The next main theorems of this paper deal with the concept of universal sets. Let $X\subseteq\mathbb{R}^{2}$, and let $\mathcal{C}$ be a class of subsets of $\mathbb{R}^{2}$ e.g. Borel sets, analytic sets, or weakly regular sets. We say that $X$ is weakly universal for pinned distances for $\mathcal{C}$ if, for every $Y\in\mathcal{C}$
(3) $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))=\min\{1,\dim_{H}(Y)\}.$
We can strengthen this notion by requiring the existence of specific points in $X$ achieving the desired bound. Hence, we say that $X$ is universal for pinned distances for $\mathcal{C}$ if, for every $Y\in\mathcal{C}$ there exists some $x\in X$ such that
(4) $\dim_{H}(\Delta_{x}(Y))=\min\{\dim_{H}(Y),1\}$
It is not obvious that, for a non-trivial class $\mathcal{C}$, small universal sets exist. The final section of this paper establishes the existence of such sets. For weakly universal sets, we have the following:
Theorem 4.
Let $X\subseteq\mathbb{R}^{2}$ be a Borel set with $\dim_{H}(X)=\dim_{P}(X)$. Then $X$ is weakly universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$ satisfying $\dim_{H}(Y)<\dim_{H}(X)$.
Unpacking the definitions, this theorem shows that there is a fixed set $X\subseteq\mathbb{R}^{2}$ such that, for every $Y$ with $\dim_{H}(Y)<\dim_{H}(X)$, $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))$ is maximal. Moreover, combined with a previous theorem of the authors [8], we have the following corollary.
Corollary 5.
Let $X\subseteq\mathbb{R}^{2}$ be a Borel set with $\dim_{H}(X)=\dim_{P}(X)>1$. Then $X$ is weakly universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$.
We are able to improve upon this corollary if we further require that $X$ is compact and AD-regular.
Theorem 6.
Let $X\subseteq\mathbb{R}^{2}$ be a compact, Ahlfors-David regular set such that $\dim_{H}(X)>1$. Then $X$ is universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$.
In particular, if we take any self-similar four-corner Cantor set $C$ of dimension strictly greater than $1$, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in C$ such that
$\dim_{H}(\Delta_{x}Y)=\min\{\dim_{H}(Y),1\}.$
2. Preliminaries
2.1. Kolmogorov complexity and effective dimension
We use the algorithmic, or effective, techniques established in [13, 1]. For a more detailed overview of these ideas, and their uses in (classical) geometric measure theory, see, e.g., [15, 16, 17, 18, 26].
The Kolmogorov complexity of a point $x\in\mathbb{R}^{m}$ at precision $r\in\mathbb{N}$ is the length of the shortest program $\pi$ that outputs a precision-$r$ rational estimate for $x$. Formally, this is
$K_{r}(x)=\min\left\{K(p)\,:\,p\in B_{2^{-r}}(x)\cap\mathbb{Q}^{m}\right\}\,,$
where $B_{\varepsilon}(x)$ denotes the open ball of radius $\varepsilon$ centered on $x$. Note that this implies that the Kolmogorov complexity of a point is non-decreasing in precision. The conditional Kolmogorov complexity of $x$ at precision $r$ given $y\in\mathbb{R}^{n}$ at precision $s\in\mathbb{R}^{n}$ is
$K_{r,s}(x\mid y)=\max\big{\{}\min\{K_{r}(p\mid q)\,:\,p\in B_{2^{-r}}(x)\cap\mathbb{Q}^{m}\}\,:\,q\in B_{2^{-s}}(y)\cap\mathbb{Q}^{n}\big{\}}\,.$
When the precisions $r$ and $s$ are equal, we abbreviate $K_{r,r}(x\mid y)$ by $K_{r}(x\mid y)$. As a matter of notational convenience, if we are given a non-integral positive real as a precision parameter, we will always round up to the next integer.
A basic property, proven by Case and J. Lutz [2] shows that the growth rate of the Kolmogorov complexity of a point is essentially bounded from above by the dimension of the ambient space. Applying this to $\mathbb{R}^{2}$ $\mathbb{R}^{2}$, for any $\varepsilon>0$, for sufficiently large $s$ we have that
$K_{r+s}(x)\leq K_{r}(x)+2s+\varepsilon s$
We may relativize the definitions in this section to an arbitrary oracle set $A\subseteq\mathbb{N}$. We will frequently consider the complexity of a point $x\in\mathbb{R}^{n}$ relative to a point $y\in\mathbb{R}^{m}$, i.e., relative to an oracle set $A_{y}$ that encodes the binary expansion of $y$ is a standard way. We then write $K^{y}_{r}(x)$ for $K^{A_{y}}_{r}(x)$. Oracle access to the entire binary expansion of a point is no less useful than conditional access to that binary expansion only up to a certain precision. Thus, we note that, for every $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$,
(5) $K_{s,r}(x\mid y)\geq K^{y}_{s}(x)-O(\log r)-O(\log s),$
for every $s,r\in\mathbb{N}$
One of the most useful properties of Kolmogorov complexity is that it obeys the symmetry of information. Lemma 7, an expression of symmetry of information for points in $\mathbb{R}^{n}$, was proved in [18].
Lemma 7 ([18]).
For every $m,n\in\mathbb{N}$, $x\in\mathbb{R}^{m}$, $y\in\mathbb{R}^{n}$, and $r,s\in\mathbb{N}$ with $r\geq s$,
  1. (i)
    $|K_{r}(x\mid y)+K_{r}(y)-K_{r}(x,y)\big{|}\leq O(\log r)+O(\log\log|y|)\,.$
  2. (ii)
    $|K_{r,s}(x\mid x)+K_{s}(x)-K_{r}(x)|\leq O(\log r)+O(\log\log|x|)\,.$
A consequence of Lemma 7, is the following.
Lemma 8 ([18]).
Let $m,n\in\mathbb{N}$, $x\in\mathbb{R}^{m}$, $z\in\mathbb{R}^{n}$, $\varepsilon>0$ and $r\in\mathbb{N}$. If $K^{x}_{r}(z)\geq K_{r}(z)-O(\log r)$, then the following hold for all $s\leq r$.
  1. (i)
    $K^{x}_{s}(z)\geq K_{s}(z)-O(\log r)\,.$
  2. (ii)
    $K_{s,r}(x\mid z)\geq K_{s}(x)-O(\log r)\,.$
In this paper, we use Mayordomo’s characterization [20] of the effective Hausdorff dimension of a point $x\in\mathbb{R}^{n}$. This is defined as
$\dim(x)=\liminf_{r\to\infty}\frac{K_{r}(x)}{r}\,.$
The effective packing dimension of a point $x\in\mathbb{R}^{n}$ is
$\operatorname{Dim}(x)=\limsup_{r\to\infty}\frac{K_{r}(x)}{r}\,.$
We can relativize both definitions, so that the effective Hausdorff and packing dimension with respect to an oracle $A\subseteq\mathbb{N}$ are
$\dim^{A}(x)=\liminf\limits_{r\to\infty}\frac{K^{A}_{r}(x)}{r}$ and $\operatorname{Dim}^{A}(x)=\limsup\limits_{r\to\infty}\frac{K^{A}_{r}(x)}{r}$
The following lemma expresses that we can often use an oracle to lower the complexity of points at specified precisions without unduly affecting the complexity of other objects.
Lemma 9 ([18]).
Let $z\in\mathbb{R}^{2}$, $\eta\in\mathbb{Q}_{+}$, and $r\in\mathbb{N}$. Then there is an oracle $D=D(r,z,\eta)$ with the following properties.
  • (i)
    For every $t\leq r$,
    $K^{D}_{t}(z)=\min\{\eta r,K_{t}(z)\}+O(\log r)\,.$
  • (ii)
    For every $m,t\in\mathbb{N}$ and $y\in\mathbb{R}^{m}$,
    $K^{D}_{t,r}(y\mid z)=K_{t,r}(y\mid z)+O(\log r)\,,$
    and
    $K_{t}^{z,D}(y)=K_{t}^{z}(y)+O(\log r)\,.$
  • (iii)
    If $B\subseteq\mathbb{N}$ satisfies $K^{B}_{r}(z)\geq K_{r}(z)-O(\log r)$, then
    $K_{r}^{B,D}(z)\geq K_{r}^{D}(z)-O(\log r)\,.$
  • (iv)
    For every $t\in\mathbb{N}$, $u\in\mathbb{R}^{n},w\in\mathbb{R}^{m}$
    $K_{r,t}(u\mid w)\leq K^{D}_{r,t}(u\mid w)+K_{r}(z)-\eta r+O(\log r)\,.$
In particular, this oracle $D$ encodes $\sigma$, the lexicographically first time-minimizing witness to $K(z{\upharpoonright}r\mid z{\upharpoonright}s)$, where $s=\max\{t\leq r\,:\,K_{t-1}(z)\leq\eta r\}$.
2.2. The Point-to-Set Principle
The point-to-set principle shows that the Hausdorff and packing dimension of a set can be characterized by the effective Hausdorff and effective packing dimension of its individual points. It is the bridge that we use to relate our effective results to our classical theorems in geometric measure theory.
Theorem 10 (Point-to-set principle [14]).
Let $n\in\mathbb{N}$ and $E\subseteq\mathbb{R}^{n}$. Then
$\operatorname{dim_{H}}(E)=\adjustlimits{\min}_{A\subseteq\mathbb{N}}{\sup}_{x\in E}\dim^{A}(x).$
$\operatorname{dim_{P}}(E)=\adjustlimits{\min}_{A\subseteq\mathbb{N}}{\sup}_{x\in E}\operatorname{Dim}^{A}(x).$
The point-to-set principle is extremely useful, but for some applications, we will also need to be able to say something more specific about which oracles are Hausdorff oracles for a given set. Hitchcock [10] and J. Lutz [13] showed that this is possible for a restricted class of sets.
A set $E\subseteq\mathbb{R}^{n}$ is effectively compact relative to $A$ if the set of finite open covers of $E$ by rational balls is computably enumerable relative to $A$. We will use the fact that every compact set is effectively compact relative to some oracle. Hitchcock [10] and J. Lutz [13] proved that, if $E$ is effectively compact relative to an oracle $A$, then $A$ is a Hausdorff oracle for $E$ (i.e., $A$ testifies to the point-to-set principle).
3. Projection theorem
The purpose of this section is to prove the following (algorithmic) theorem.
Theorem 11.
Let $x\in\mathbb{R}^{2},e\in S^{1},\sigma\in\mathbb{Q}\cap(0,1),\varepsilon>0,C\in\mathbb{N},A\subseteq\mathbb{N}$, and $t<r\in\mathbb{N}$. Let $D\geq\operatorname{Dim}^{A}(x)$ and $\alpha=\min\{D,1+\sigma\}$. Suppose that $r$ is sufficiently large and that the following hold
  1. (1)
    $0<\sigma<\dim^{A}(x)\leq 1$,
  2. (2)
    $t\geq\max\{\frac{\sigma(\sigma+1-\alpha)}{\alpha+\sigma+\sigma^{2}-\sigma\alpha}r,\frac{r}{C}\}$,
  3. (3)
    $K_{s}^{A,x}(e)\geq\sigma s-C\log s$ for all $s\leq t$
Then
$K_{r}^{A}(x|p_{e}x,e)\leq K^{A}_{r}(x)-\sigma r+\frac{(\alpha-\sigma)\sigma}{\alpha}(r-t)+\varepsilon r.$
This theorem gives strong upper bounds on the number of bits needed to compute a point $x\in\mathbb{R}^{2}$ to precision $r$, if the direction $e$ and the projection $p_{e}x$ are known to precision $r$.
The proof of this theorem involves several steps. The first, detailed in Section 3.1, gives tight bounds on the complexity $K^{A}_{b,b,b,a}(x\mid p_{e}x,e,x)$ for special intervals of precisions $[a,b]$. The intervals of precisions considered depend on the behavior of the complexity function $K^{A}(x)$ on the interval $[a,b]$.
The auxiliary results of Section 3.1 are only applicable on a given interval of precisions when the complexity function $K_{s}^{A}(x)$ is of a certain form on that interval. In Section 3.2 we construct a partition of $[1,r]$ such that every interval in the partition is of the necessary type to apply the results of Section 3.1.
Finally, in Section 3.3 we use this partition to prove our main theorem.
3.1. Projection lemmas
To begin, we recall two lemmas which we will need for the results of this section.
Lemma 12 ([18]).
Let $A\subseteq\mathbb{N}$, $x\in\mathbb{R}^{2}$, $e\in S^{1}$, and $r\in\mathbb{N}$. Let $w\in\mathbb{R}^{2}$ such that $p_{e}x=p_{e}w$ up to precision $r$. Then
$K^{A}_{r}(w)\geq K^{A}_{t}(x)+K^{A}_{r-t,r}(e\mid x)+O(\log r)\,,$
where $t:=-\log|x-w|$.
The next lemma was essentially proved by Lutz and Stull [18], with slight modifications for the purposes of this paper.
Lemma 13.
Suppose that $x\in\mathbb{R}^{2}$, $e\in S^{1}$, $t<r\in\mathbb{N}$, $\sigma\in(0,1]$, $B\subseteq\mathbb{N}$, and $\eta,\varepsilon\in\mathbb{Q}^{+}$ satisfy the following conditions
  1. (1)
    $K^{B}_{r}(x)\leq\eta r+\frac{\varepsilon}{2}r$
  2. (2)
    For every $w\in B_{2^{-t}}(x)$ such that $p_{e}w=p_{e}x$,
    $K^{B}_{r}(w)\geq\eta r+\min\{\varepsilon r,\sigma(r-s)-\varepsilon r\}$
    where $s$ = $-\log|x-w|\leq r$
Then,
$K^{B}_{r,r,r,t}(x|p_{e}x,e,x)\leq\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+O(\log r)$
Equipped with these facts, we can prove the following two lemmas which bound $K_{r,r,r,t}(x|p_{e}x,e,x)$ on certain kinds of intervals.
Lemma 14.
Let $x\in\mathbb{R}^{2}$, $e\in S^{1}$, $\varepsilon\in\mathbb{Q}^{+}$, $C>0$, $\sigma\in(0,1]\cap\mathbb{Q}_{+}$, $A\subseteq{N}$, and $t<r\in\mathbb{N}$ be given. For sufficiently large $r$, if
  1. (1)
    $K_{s}^{A,x}(e)\geq\sigma s-C\log r$ for all $s\leq r-t$, and
  2. (2)
    $K^{A}_{s,t}(x)\geq\sigma(s-t)-\varepsilon r$ for all $t\leq s\leq r$.
Then,
$K^{A}_{r,r,r,t}(x|p_{e}x,e,x)\leq K^{A}_{r,t}(x|x)-\sigma(r-t)+\frac{7\varepsilon}{\sigma}r$
Proof.
Let $\eta\in\mathbb{Q}_{+}$ be such that $\eta r=K^{A}_{t}(x)+\sigma(r-t)-3\varepsilon r$. Note that $K(\eta\mid\varepsilon,\sigma)=O(\log r)$. In particular, since $\sigma$ and $\varepsilon$ are fixed rationals and $r$ is sufficiently large, $K(\eta,\varepsilon)=O(\log r)$. Let $D=D^{A}(r,x,\eta)$ be the oracle of Lemma 9, relative to $A$. Then, $K_{r}^{A,D}(x)\leq(\eta+\frac{\varepsilon}{2})r$, so the first condition of Lemma 13 is satisfied relative to $(A,D)$. It remains to verify that the second condition of Lemma 13 holds. Let $w\in\mathbb{R}^{2}$ such that $p_{e}x=p_{e}w$ and $t\leq-\log|x-w|=s\leq r$. Then, by Lemma 12,
$K_{r}^{A,D}(w)$ $\geq K_{s}^{A,D}(x)+K_{r-s,r}^{A,D}(e|x)-O(\log r)$
$\geq K_{s}^{A,D}(x)+K^{A}_{r-s,r}(e|x)-O(\log r)$
$\geq K_{s}^{A,D}(x)+K^{A,x}_{r-s,r}(e)-O(\log r)$
$\geq K_{s}^{A,D}(x)+\sigma(r-s)-O(\log r)$
By the properties of $D$, there are two cases: either $K_{s}^{A,D}(x)=K^{A}_{s}(x)+O(\log r)$ or $K_{s}^{A,D}(x)=\eta r+O(\log r)$. In the first case, we have
$K_{r}^{A,D}(w)$ $\geq K^{A}_{s}(x)+\sigma(r-s)-O(\log r)$
$\geq K^{A}_{t}(x)+K^{A}_{s,t}(x|x)+\sigma(r-s)-O(\log r)$
$\geq K^{A}_{t}(x)+\sigma(s-t)+\sigma(r-s)-\varepsilon r-O(\log r)$
$=K^{A}_{t}(x)+\sigma(r-t)-\varepsilon r-O(\log r)$
$=(\eta+2\varepsilon)r-O(\log r)$
$\geq(\eta+\varepsilon)r$
In the second case, we have
$K_{r}^{A,D}(w)$ $\geq\eta r+\sigma(r-s)-O(\log r)$
$\geq\eta r+\sigma(r-s)-\varepsilon r$
Therefore, the second condition of Lemma 13 is satisfied, and so we conclude
$K^{A,D}_{r,r,r,t}(x|p_{e}x,e,x)\leq\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+O(\log r)$
Then by Lemma 9, and the symmetry of information, we have
$K^{A}_{r,r,r,t}(x|p_{e}x,e,x)$ $\leq K^{A,D}_{r,r,r,t}(x|p_{e}x,e,x)+K_{r}^{A}(x)-\eta r+O(\log r)$
$\leq\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+K_{r}^{A}(x)-\eta r+O(\log r)$
$\leq K^{A}_{r,t}(x\mid x)-\sigma(r-t)+3\varepsilon r+\frac{3\varepsilon}{\sigma}r+O(\log r)$
$\leq K_{r,t}^{A}(x|x)-\sigma(r-t)+\frac{7\varepsilon}{\sigma}r$
Now, we state the corresponding lemma for the second kind of interval, and prove it (in an almost identical manner).
Lemma 15.
Let $x\in\mathbb{R}^{2}$, $e\in S^{1}$, $\varepsilon\in\mathbb{Q}^{+}$, $C\in\mathbb{N}$, $\sigma\in(0,1]\cap\mathbb{Q}_{+}$, $A\subseteq\mathbb{N}$, and $t<r\in\mathbb{N}$ be given. For sufficiently large $r$, if
  1. (1)
    $K_{s}^{A,x}(e)\geq\sigma s-C\log r$ for every $s\leq r-t$, and
  2. (2)
    $K^{A}_{r,s}(x)\leq\sigma(r-s)+\varepsilon r$, for every $t\leq s\leq r$.
Then,
$K^{A}_{r,r,r,t}(x|p_{e}x,e,x)\leq\frac{7\varepsilon}{\sigma}r$
Proof.
Pick $\eta\in\mathbb{Q}_{+}$ be such that $\eta r=K^{A}_{r}(x)-3\varepsilon r$. We again note that $K(\eta,\varepsilon)=O(\log r)$. Let $D=D^{A}(r,x,\eta)$ be the oracle of Lemma 9, relative to $A$. Then, $K_{r}^{A,D}(x)\leq(\eta+\frac{\varepsilon}{2})r$, so the first condition of Lemma 13 is satisfied relative to $(A,D)$. It remains to verify the second. Let $w\in\mathbb{R}^{2}$ such that $p_{e}x=p_{e}w$ and $t\leq-\log|x-w|=s\leq r$. Then, by Lemma 12,
$K_{r}^{A,D}(w)$ $\geq K_{s}^{A,D}(x)+K_{r-s,r}^{A,D}(e|x)-O(\log r)$
$\geq K_{s}^{A,D}(x)+\sigma(r-s)-O(\log r)$
By the properties of $D$, there are two cases: either $K_{s}^{A,D}(x)=K^{A}_{s}(x)+O(\log r)$ or $K_{s}^{A,D}(x)=\eta r+O(\log r)$. In the first case, we have
$K_{r}^{A,D}(w)$ $\geq K^{A}_{s}(x)+\sigma(r-s)-O(\log r)$
$\geq K^{A}_{s}(x)+K^{A}_{r,s}(x|x)-\varepsilon r-O(\log r)$
$=K^{A}_{r}(x)-\varepsilon r-O(\log r)$
$=(\eta+2\varepsilon)r-O(\log r)$
$\geq(\eta+\varepsilon)r$
In the second case, we have
$K_{r}^{A,D}(w)$ $\geq\eta r+\sigma(r-s)-O(\log r)$
$\geq\eta r+\sigma(r-s)-\varepsilon r$
Therefore, the second condition of Lemma 13 is satisfied, and so
$K^{A,D}_{r,r,r,t}(x|p_{e}x,e,x)\leq\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+O(\log r)$
Then by Lemma 9, and the symmetry of information, we have
$K^{A}_{r,r,r,t}(x|p_{e}x,e,x)$ $\leq K^{A,D}_{r,r,r,t}(x|p_{e}x,e,x)+K_{r}^{A}(x)-\eta r+O(\log r)$
$\leq\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+K_{r}^{A}(x)-\eta r+O(\log r)$
$\leq 3\varepsilon r+\frac{3\varepsilon}{\sigma}r+O(\log r)$
$\leq\frac{7\varepsilon}{\sigma}r.$
In addition to the previous lemmas, there is another bound which can be useful. In the setting of [8] and [27], since $\sigma=1$, the bound
$K^{A}_{r,r,r,t}(x\mid p_{e}x,e,x)\leq\max\{K_{r,t}^{A}(x)-\sigma(r-t),0\}\leq r-t$
held automatically on appropriate intervals. We can recover the same bound in the setting $\sigma<1$ with the following lemma.
Lemma 16.
Let $A\subseteq\mathbb{N}$, $x\in\mathbb{R}^{2},e\in\mathcal{S}^{1}$, and $a\leq b\in\mathbb{N}$. If $b$ is sufficiently large, then
$K^{A}_{b,b,b,a}(x\mid p_{e}x,e,x)\leq b-a+O(\log b)$.
Proof.
The proof follows from three simple observations:
  • $e^{\perp}$ is easily computable to precision $b$ given $e$ to precision $b$,
  • $x=(p_{e}x)e+(p_{e^{\perp}}x)e^{\perp}$ is easily computable to precision $b$ given these four objects to precision $b$.
  • $p_{e^{\perp}}x$ is easily computable to precision $a<b$ given $e^{\perp}$ to precision $b$ and $x$ to precision $a$
Applying these observations in order, have
$K^{A}_{b,b,b,a}(x\mid p_{e}x,e,x)$ $\leq K^{A}_{b,b,b,b,a}(x\mid p_{e}x,e,e^{\perp},x)+O(\log b)$
$\leq K^{A}_{b,b,a}((p_{e^{\perp}}x)\mid e^{\perp},x)+O(\log b)$
$\leq K^{A}_{b,a}(p_{e^{\perp}}x\mid p_{e^{\perp}}x)+O(\log b).$
Since $p_{e^{\perp}}x$ is a real number, the conclusion follows.
3.2. Admissible partitions
For the remainder of this section, we fix a point $x\in\mathbb{R}^{2}$ and oracle $A\subseteq\mathbb{N}$. First, after [27], we extend the complexity function $K_{s}^{A}(x)$ to real numbers via $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$, the piece-wise linear function such that $f(x)=K^{A}_{s}(x)$ on the integers such that
$f(a)=f(\lfloor a\rfloor)+(a-\lfloor a\rfloor)(f(\lceil a\rceil)-f(\lfloor a\rfloor))$,
for any non-integer $a$. Now, we can introduce several kinds of intervals corresponding to or identical to those in [27].
  • An interval $[a,b]$ is $\boldsymbol{\sigma}$-teal if for every $s\in[a,b]$, $f(b)-f(s)\leq\sigma(b-s)$.
  • An interval $[a,b]$ is $\boldsymbol{\sigma}$-yellow if for every $s\in[a,b]$, $f(s)-f(a)\geq\sigma(s-a)$.
  • An interval $[a,b]$ is $\boldsymbol{\sigma}$-green if it is both $\sigma$-teal and $\sigma$-yellow, and $b-a\leq t$.
  • An interval $[a,b]$ is red if $f(d)-f(c)>d-c$ for every $a\leq c<d\leq b$.
  • An interval $[a,b]$ is blue if $f$ is constant on $[a,b]$.
Intervals of these types have desirable properties. In particular, using Lemmas 14, 15 and 16 we are able to give tight bounds on the complexity $K^{A}(x\mid p_{e}x,e,x)$.
Proposition 17.
Let $A\subseteq\mathbb{N}$, $x\in\mathbb{R}^{2},e\in\mathcal{S}^{1},\sigma\in(0,1]\cap\mathbb{Q}_{+},\varepsilon>0$, and $a<b\in\mathbb{R}_{+}$. Suppose that $b$ is sufficiently large and $K^{A,x}_{t}(e)\geq\sigma t-O(\log b)$, for all $t\leq b-a$. Then the following hold.
  1. (1)
    If $[a,b]$ is $\sigma$-yellow,
    $K^{A}_{b,b,b,a}(x\mid p_{e}x,e,x)\leq\min\{K^{A}_{b,a}(x\mid x)-\sigma(b-a),b-a\}+\varepsilon b$.
  2. (2)
    If $[a,b]$ is $\sigma$-teal,
    $K^{A}_{b,b,b,a}(x\mid p_{e}x,e,x)\leq\varepsilon b$.
Proof.
The claim follows from Lemmas 15, 14 and 16. ∎
Our goal is to partition the interval $[1,r]$ into teal and yellow intervals, and sum the contributions of each using the symmetry of information. To accomplish all of this, we define an $(M,r,t,\sigma)$-admissible partition to be a partition $[a_{i},a_{i+1}]_{i=0}^{k}$ of $[1,r]$ such that:
  1. (1)
    $k<M$
  2. (2)
    Every interval in the partition is either $\sigma$-yellow or $\sigma$-teal (or both).
  3. (3)
    No interval has length more than $t$.
Lemma 18.
Suppose that $A\subseteq\mathbb{N}$, $x\in\mathbb{R}^{2}$, $e\in\mathcal{S}^{1}$, $\sigma\in(0,1]\cap\mathbb{Q}_{+}$, $\varepsilon>0$, $C\in\mathbb{N}$, $t,r\in\mathbb{N}$ satisfy
  1. (1)
    $\dim^{A}(x)>\sigma$,
  2. (2)
    $K^{A,x}_{s}(e)\geq\sigma s-C\log s$ for all $s\leq t$.
If $\mathcal{P}=\{[a_{i},a_{i+1}]\}_{i=0}^{k}$ is an $(10C,r,t,\sigma)$-admissible partition, and $r$ is sufficiently large, then
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq\varepsilon r+\sum\limits_{i\in\textbf{Bad}}\min\{K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\},$
where
Bad =$=\{i\leq k\mid[a_{i},a_{i+1}]\notin T_{\sigma}\}$.
Proof.
We assume $r$ is large enough so that, for any $b>\log r$, the conditions of Proposition 17 hold for $\frac{\varepsilon}{11C}$. Starting with $\mathcal{P}$ a $(10C,r,t,\sigma)$-admissible partition, suppose $j$ is the largest integer such that $a_{j}<\log r$. Remove the intervals $[1,a_{1}],...[a_{j-1},a_{j}]$ from $\mathcal{P}$ and add $[1,a_{j}]$, still calling the modified partition $\mathcal{P}$ and re-indexing it for ease of notation. Applying Proposition 17 we see that on the teal intervals of $\mathcal{P}$,
$K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)\leq\frac{\varepsilon}{11C}r,$
and on the yellow intervals
$K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)\leq\min\{K^{A}_{a_{i+1}a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\}+\frac{\varepsilon}{11C}r\$
Applying symmetry of information no more than $10C$ times, we obtain
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq\sum\limits_{\mathcal{P}}K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)+O(\log r)$
$\leq\sum\limits_{\mathcal{P}\setminus[1,a_{1}]}K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)+O(\log r)$
$\leq\sum\limits_{i\in\textbf{Teal}}K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)$
$\qquad+\sum\limits_{i\in\textbf{Bad}}K^{A}_{a_{i+1},a_{i+1},a_{i+1},a_{i}}(x\mid p_{e}x,e,x)+O(\log r)$
$\leq\sum\limits_{i\in\textbf{Bad}}\min\{K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\}+\frac{\varepsilon}{11C}$
$\qquad+\sum\limits_{i\in\textbf{Teal}}\frac{\varepsilon}{11C}r+O(\log r)$
$\leq\sum\limits_{i\in\textbf{Bad}}\min\{K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\}+\frac{10\varepsilon}{11}r+O(\log r).$
$\leq\sum\limits_{i\in\textbf{Bad}}\min\{K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\}+\varepsilon r.$
An immediate consequence of the above is that, if $B$ is the total length of the bad intervals,
$K^{A}_{r}(x\mid p_{e}x,e)\leq B+\varepsilon r$
In the proof of Theorem 31, we will need a special partition. Specifically, we need a partition which will maximize the proportion of $[1,r]$ covered by of $\sigma$-green intervals.
Lemma 19.
Let $x\in\mathbb{R}^{2},A\subseteq\mathbb{N}$, $\sigma\in(0,1]\cap\mathbb{Q}$ and $t,r\in\mathbb{N}$ be given. Then there exists a partition $\mathcal{P}$ of $[1,r]$ with the following properties:
  • Every interval in $\mathcal{P}$ is red, $\sigma$-green, or blue
  • $\mathcal{P}$ maximizes the amount of $\sigma$-green, in the sense that, if $s\in[1,r]$ is contained in a $\sigma$-green interval, then $s$ is contained in a $\sigma$-green interval in $\mathcal{P}$.
  • If $[r_{i},r_{i+1}]$ is red, $[r_{i+1},r_{i+2}]$ cannot be blue.
  • The green interval(s) in every red–green–blue sequence of $\mathcal{P}$ have (total) length at least $t$.
Proof.
The construction of such a partition is essentially the same as in [8]; for completeness, we include the details.
Given $x$, $A$, $t<r$, and $\sigma$, consider the set of all $\sigma$-green intervals. First, we refine this collection via the following greedy strategy.
  1. (1)
    Remove any $\sigma$-green interval strictly contained in another from the collection.
  2. (2)
    Moving from left to right, replace the first overlapping $\sigma$-green intervals $[a,b]$ and $[c,d]$ with $[a,c]$ and $[b,d]$, noting that $[b,d]$ is $\sigma$-green.
  3. (3)
    Remove $[b,d]$ if it is now strictly contained in some interval in the remaining collection
  4. (4)
    Check whether any overlapping $\sigma$-green intervals remain, and if so, repeat steps 2 and 3. Otherwise, label the remaining collection $G$ and stop.
The elements of $G$ have disjoint interiors, and can be added to $\mathcal{P}$. It is clear that the second property is satisfied.
To finish the proof, observe that a blue interval $[b,c]$ following a red interval $[a,b]$ implies that $b$ is contained in a $\sigma$-green interval of length at most $t$. Thus, there are only three cases for the remaining gaps in $[1,r]\setminus G$: a gap is red, a gap is blue, or a gap is the union of a blue interval and then a red interval. In each case, simply add the correct color of interval(s) to $\mathcal{P}$ to complete the partition. Now, $\mathcal{P}$ consists of only red, $\sigma$-green, and blue intervals, so the first property is satisfied, and the above observation combined with the first two properties gives the third.
We now observe that the final property is satisfied by $\mathcal{P}$. We may reduce to the case of one $\sigma$-green interval in the sequence: there cannot be two adjacent $\sigma$-green intervals with total length less than $t$. So, it suffices to consider the red–$\sigma$-green–blue sequence $[a,b],[b,c],[c,d]$. Assume $c-b<t$. We know that $\frac{f(c)-f(b)}{c-b}=\sigma$. For all sufficiently small $\delta_{1},\delta_{2}$, $f(b-\delta_{1})<f(b)-\delta_{1}$ and $f(c+\delta_{2})=f(c)$. It follows that there is a $\sigma$-green interval $[a^{\prime},c^{\prime}]$ such that $a^{\prime}<b$ and $c^{\prime}>c$, contradicting the construction of $\mathcal{P}$. ∎
3.3. Main projection theorem
We now use the partition construction of Lemma 19 to prove our main (effective) projection theorem. The proof follows from a careful analysis of this partition.
Theorem 11.
Let $x\in\mathbb{R}^{2},e\in S^{1},\sigma\in\mathbb{Q}\cap(0,1),\varepsilon>0,C\in\mathbb{N},A\subseteq\mathbb{N}$, and $t<r\in\mathbb{N}$. Let $D\geq\operatorname{Dim}^{A}(x)$ and $\alpha=\min\{D,1+\sigma\}$. Suppose that $r$ is sufficiently large and that the following hold
  1. (1)
    $0<\sigma<\dim^{A}(x)\leq 1$,
  2. (2)
    $t\geq\max\{\frac{\sigma(\sigma+1-\alpha)}{\alpha+\sigma+\sigma^{2}-\sigma\alpha}r,\frac{r}{C}\}$,
  3. (3)
    $K_{s}^{A,x}(e)\geq\sigma s-C\log s$ for all $s\leq t$
Then
$K_{r}^{A}(x|p_{e}x,e)\leq K^{A}_{r}(x)-\sigma r+\frac{(\alpha-\sigma)\sigma}{\alpha}(r-t)+\varepsilon r.$
Proof.
First, let $\mathcal{P}$ be the partition of Lemma 19. Using a greedy strategy, we can partition the red and blue intervals in $\mathcal{P}$ into $\sigma$-yellow and $\sigma$-teal intervals of length at most $t$. Abusing notation, we call this partition $\mathcal{P}$. Let $S$ be the number of $\sigma$-RGB sequences in this partition. We prove the theorem by cases.
Case $S=0$: In this case, $\mathcal{P}$ is almost entirely red and $\sigma$-green. In particular, there is some constant $c$ such that for any real $s>c$, $s$ is not contained in a blue interval of $\mathcal{P}$. A simple greedy construction gives an admissible partition of $[c,r]$ consisting entirely of $\sigma$-yellow intervals (which we will refer to as $\mathcal{P}$). Then, by Lemma 18,
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq\frac{\varepsilon}{2}r+\sum\limits_{i\in\textbf{Bad}}K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i})$
$\leq(\alpha-\sigma)r+\varepsilon r.$
Case $S=1$: Suppose the green portion of our red-green-blue sequence starts at precision $r_{1}$ and ends at precision $r_{1}+s$, for some $s\geq t$. Let $r_{2}\geq r_{1}+s$ be the minimal precision such that $[r_{2},r]$ is the union of yellow intervals [^2] No such precision necessarily exists. In this case, set $r_{2}=r$..
We first assume that $B>\frac{\sigma}{\alpha}(r-t)$, where $B$ is the total length of the non-teal intervals in $\mathcal{P}$. Then, by Lemma 18, we have
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq K^{A}_{r}(x)-\sigma s-\sigma B+\varepsilon r$
$<K^{A}_{r}(x)-\sigma t-\sigma\frac{\sigma}{\alpha}(r-t)+\varepsilon r$
$=K^{A}_{r}(x)-\sigma r+\frac{(\alpha-\sigma)\sigma}{\alpha}(r-t)+\varepsilon r,$
as required. Now suppose that $B\leq\frac{\sigma}{\alpha}(r-t)$. Note that this immediately implies that $r_{1}\leq\frac{\sigma}{\alpha}(r-t)$. Using the same argument as the $S=0$ case for the interval $[1,r_{1}]$ we conclude that
(6) $K^{A}_{r_{1}}(x\mid p_{e}x,e)\leq\left(\alpha-\sigma+\frac{\varepsilon}{3}\right)r_{1}$
Using the same argument for the interval $[r_{2},r]$, we have
(7) $K^{A}_{r,r,r,r_{2}}(x\mid p_{e}x,e,x)\leq K^{A}_{r,r_{2}}(x\mid x)-\sigma(r-r_{2})+\frac{\varepsilon}{3}r.$
Finally, since the interval $[r_{1},r_{2}]$ is $\sigma$-green, we have
(8) $K^{A}_{r_{2},r_{2},r_{2},r_{1}}(x\mid p_{e}x,e,x)\leq\frac{\varepsilon}{3}r.$
Combining inequalities (6), (7) and (8), the bound on $r_{1}$, and symmetry of information, yields
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq\left(\alpha-\sigma\right)r_{1}+K^{A}_{r,r_{2}}(x\mid x)-\sigma(r-r_{2})+\varepsilon r$
$\leq\left(\alpha-\sigma\right)r_{1}+K^{A}_{r}(x)-\sigma r+\varepsilon r$
$\leq\left(\alpha-\sigma\right)\frac{\sigma}{\alpha}(r-t)+K^{A}_{r}(x)-\sigma r+\varepsilon r,$
and the proof for the $S=1$ case is complete.
Case $S\geq 2$: Let $L$ be the total length of the $\sigma$-green intervals, and $B$ be the total length of the non-teal intervals. Recall that, by Lemma 18, $K^{A}_{r}(x\mid p_{e}x,e)\leq B$. We also have that
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq K^{A}_{r}(x)-\sigma L-\sigma B+\varepsilon r$
$\leq K^{A}_{r}(x)-2\sigma t-\sigma B+\varepsilon r.$
Hence we see that
(9) $K^{A}_{r}(x\mid p_{e}x,e)\leq\min\{K^{A}_{r}(x)-\sigma B-2\sigma t,B\}+\varepsilon r.$
By the above inequality, if
$B\leq K^{A}_{r}(x)-\sigma r+\frac{(\alpha-\sigma)\sigma}{\alpha}(r-t)$
the proof is complete, so we assume otherwise. In this case, using (9), we see that
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq K^{A}_{r}(x)-2\sigma t-\sigma B+\varepsilon r$
$\leq(1-\sigma)K^{A}_{r}(x)+\frac{\sigma^{3}}{\alpha}r-\frac{2\alpha-(\alpha-\sigma)\sigma}{\alpha}\sigma t+\varepsilon r.$
Therefore, it suffices to show that
(10) $\frac{\sigma(1+\sigma)}{\alpha}r-\frac{\alpha-\alpha\sigma+\sigma^{2}+\sigma}{\alpha}t\leq K^{A}_{r}(x).$
Since $K^{A}_{r}(x)\geq\sigma r$, it suffices to show that
(11) $\frac{\sigma(1+\sigma)}{\alpha}r-\frac{\alpha-\alpha\sigma+\sigma^{2}+\sigma}{\alpha}t\leq\sigma r.$
We first note that, if $D\geq 1+\sigma$, then this holds trivially, and the proof is complete. Otherwise, it suffices to show that
(12) $t\geq\frac{\sigma(1+\sigma-\alpha)}{\alpha+\sigma+\sigma^{2}-\sigma\alpha}r,$
which is true by our assumed lower bound on $t$. ∎
4. Effective dimension of pinned distances
In this section, we establish results on the complexity of distances. These correspond to the projection results of the previous section.
Lemma 20.
Suppose that $B\subseteq\mathbb{N}$, $x,y\in\mathbb{R}^{2}$, $t<r\in\mathbb{N}$, $\sigma\in(0,1]$ and $\eta,\varepsilon\in\mathbb{Q}_{+}$ satisfy the following conditions.
  • (i)
    $K^{B}_{r}(y)\leq\left(\eta+\frac{\varepsilon}{2}\right)r$.
  • (ii)
    For every $z\in B_{2^{-t}}(y)$ such that $|x-y|=|x-z|$,
    $K^{B}_{r}(z)\geq\eta r+\min\{\varepsilon r,\sigma(r-s)-\varepsilon r\}\,,$
    where $s=-\log|y-z|\leq r$.
Then,
$K_{r,t}^{B,x}(y\mid y)\leq K^{B,x}_{r,t}(|x-y|\mid y)+\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+O(\log r)\,.$
Proof.
The proof is almost identical to that of [27] (Lemma 12).
Define an oracle Turing machine $M$ that does the following given oracle $(B,x)$, and inputs $\pi=\pi_{1}\pi_{2}\pi_{3}\pi_{4}$ and $q\in\mathbb{Q}^{2}$ such that $U(\pi_{2})=s_{1}\in\mathbb{N}$, $U(\pi_{3})=s_{2}\in\mathbb{N}$ and $U(\pi_{4})=(\zeta,\iota)\in\mathbb{Q}^{2}$.
$M$ first uses $q$ and oracle access to $x$ to compute a dyadic rational $d$ such that $|d-|x-q||<2^{-s_{1}}$. $M$ then computes $U^{B,x}(\pi_{1},q)=p\in\mathbb{Q}$. For every program $\tau\in\{0,1\}^{*}$ with $\ell(\tau)\leq\iota s_{2}+\frac{\zeta s_{2}}{2}$, in parallel, $M$ simulates $U(\tau)$. If one of the simulations halts with output $p_{2}\in\mathbb{Q}^{2}\cap B_{2^{-s_{1}}}(q)$ such that $|d-|x-p_{2}||<2^{-s_{2}}$, then $M^{A,e}$ halts with output $p_{2}$. Let $k_{M}$ be a constant for the description of $M$.
Let $\pi_{1}$, $\pi_{2}$, $\pi_{3}$, and $\pi_{4}$ testify to $K^{B,x}_{r,t}(|x-y|\mid y)$, $K(t)$, $K(r)$, and $K(\varepsilon,\eta)$, respectively, and let $\pi=\pi_{1}\pi_{2}\pi_{3}\pi_{4}$. It can be readily verified that $M^{B,x}(\pi,q)$ eventually halts, and outputs a rational $p\in\mathbb{Q}^{2}$, whenever $q\in B_{2^{-t}}(y)$. By the definition of $M$, $p\in B_{2^{-t}}(y)$, and $||x-p_{2}|-|x-y||<2^{1-r}$. Again by the definition of $M$, $K^{A}(p)\leq(\eta+\frac{\varepsilon}{2})r$.
Let $w\in\mathbb{R}^{2}$ such that $|w-p|<|x-y|2^{1-r}$. Since $w\in\mathbb{R}^{2}$,
$K^{A}_{r}(w)\leq(\eta+\frac{\varepsilon}{2})r+2\log\lceil|x-y|\rceil+O(\log r)$.
Let $s=-\log|y-w|$. We first assume that $\varepsilon r\leq\sigma(r-s)-\varepsilon r$. Then, by condition (ii),
$\eta r+\varepsilon r\leq(\eta+\frac{\varepsilon}{2})r+2\log\lceil|x-y|\rceil+O(\log r)$,
which is a contradiction, since $r$ was chosen to be sufficiently large. Now assume that $\sigma(r-s)-\varepsilon r<\varepsilon r$. Then,
$\eta r+\sigma(r-s)-\varepsilon r\leq(\eta+\frac{\varepsilon}{2})r+2\log\lceil|x-y|\rceil+O(\log r)$,
and so $r-s\leq\frac{3\varepsilon r}{\sigma}+O(\log r)$. Hence,
$K^{A,x}_{r}(y\mid w)\leq\frac{3\varepsilon}{\sigma}r+O(\log r)$.
Therefore,
$K^{A,x}_{r,t}(y\mid y)$ $\leq|\pi|+\frac{3\varepsilon}{\sigma}r+O(\log r)$
$=K^{A,x}_{r,t}(|x-y|\mid y|)+\frac{3\varepsilon}{\sigma}r+K(\varepsilon,\eta)+O(\log r),$
and the proof is complete. ∎
For a fixed $y\in\mathbb{R}^{2}$ and oracle $A\subseteq\mathbb{N}$, we consider the the piece-wise linear function $f$ that agrees with $K^{A}_{s}(y)$ on the integers such that
$f(a)=f(\lfloor a\rfloor)+(a-\lfloor a\rfloor)(f(\lceil a\rceil)-f(\lfloor a\rfloor))$,
for any non-integer $a$. We define $\sigma$-yellow and $\sigma$-teal in the exact same way as in Section 3.2. We say that an interval $[a,b]$ is $\sigma$-green if it is both $\sigma$-yellow and $\sigma$-teal and $b\leq 2a$.
We now prove the result which bounds the complexity of distances on special (yellow and teal) intervals. This corresponds to our projection result, Proposition 17.
Lemma 21.
Let $A\subseteq\mathbb{N}$, $x,y\in\mathbb{R}^{2}$, $e=\frac{x-y}{|x-y|}$, $\sigma\in\mathbb{Q}\cap(0,1)$, $C\in\mathbb{N}$, $\varepsilon>0$, and $t<r\leq 2t$. Suppose $r$ is sufficiently large and
$K^{A}_{s,r}(e\mid y)\geq\sigma s-C\log r$, for all $s\leq r-t$.
Then the following hold.
  1. (1)
    If $[t,r]$ is $\sigma$-yellow and $t\leq r\leq 2t$,
    $K^{A,x}_{r}(y\mid|x-y|,y)\leq K^{A}_{r-t}(y\mid y)-\sigma(r-t)+\epsilon r$.
  2. (2)
    If $[t,r]$ is $\sigma$-teal and $t\leq r\leq 2t$,
    $K^{A,x}_{r}(y\mid|x-y|,y)\leq\epsilon r$.
Proof.
Let $\varepsilon^{\prime}\in\mathbb{Q}$ be sufficiently small. Let $\eta\in\mathbb{Q}$ such that
$\eta r=\begin{cases}K^{A}_{t}(y)+\sigma(r-t)-3\varepsilon^{\prime}r&\text{ if $[t,r]$ is yellow}\\ K^{A}_{r}(y)-3\varepsilon^{\prime}r&\text{ if $[t,r]$ is teal}\end{cases}$
Let $D=D^{A}(r,y,\eta)$ be the oracle of Lemma 9. Note that, by Lemma 9(i), $K^{A,D}_{r}(y)\leq\eta r+\frac{\varepsilon^{\prime}r}{2}$.
Suppose that $w\in B_{2^{-t}}(y)$ such that $|x-w|=|x-y|$. Let $s=-\log|w-y|$. Our goal is to show a sufficiently large lower bound on the complexity of $w$, so that we may apply Lemma 20. We first note that, since $y$ and $w$ agree to precision at least $t$,
$|p_{e}y-p_{e}w|<2^{-2t+c}\leq 2^{-r+c}$,
where $c=\log|x-y|$. Therefore, by Lemma 12,
$K^{A,D}_{r}(w)$ $\geq K^{A,D}_{r-c}(w)$
$\geq K^{A,D}_{s}(y)+K^{A,D}_{r-c-s,r-c}(e\mid y)-O(\log r)$
$\geq K^{A,D}_{s}(y)+K^{A,D}_{r-s,r}(e\mid y)-O(\log r)$
$\geq K^{A,D}_{s}(y)+K^{A}_{r-s,r}(e\mid y)-O(\log r)$
$\geq K^{A,D}_{s}(y)+\sigma(r-s)-\frac{\varepsilon^{\prime}r}{2}-O(\log r).$
By the properties of our oracle $D$, there are two cases to consider. If $K^{A,D}_{s}(y)=K^{A}_{s}(y)-O(\log r)$, then
(13) $K^{A,D}_{r}(w)\geq K^{A}_{s}(y)+\sigma(r-s)-\frac{\varepsilon^{\prime}r}{2}-O(\log r).$
It is not difficult to verify that, regardless of whether $[t,r]$ is yellow or teal,
(14) $K^{A,D}_{r}(w)\geq(\eta+\varepsilon^{\prime})r.$
For the other case, when $K^{A,D}_{s}(y)=\eta r-O(\log r)$, we again have
$K^{D}_{r}(w)$ $\geq\eta r+\sigma(r-s)-\frac{\varepsilon^{\prime}r}{2}-O(\log r)$
$\geq\eta r+\sigma(r-s)-\varepsilon^{\prime}r$
Thus the conditions of Lemma 20 are satisfied. Applying this, relative to $(A,D)$, yields
(15) $K_{r,t}^{A,D,x}(y\mid y)\leq K^{A,D,x}_{r,t}(|x-y|\mid y)+\frac{3\varepsilon^{\prime}}{\sigma}r+O(\log r),$
or, equivalently,
(16) $K_{r,r,t}^{A,D,x}(y\mid|x-y|,y)\leq\frac{3\varepsilon^{\prime}}{\sigma}r+O(\log r).$
Using Lemma 9(iv), we conclude that
$K_{r,r,t}^{A,x}(y\mid|x-y|,y)$ $\leq K_{r,r,t}^{A,D,x}(y\mid|x-y|,y)+K^{A}_{r}(y)-\eta r$
$\leq K^{A}_{r}(y)-\eta r+\frac{3\varepsilon^{\prime}}{\sigma}r+O(\log r),$
and the conclusion follows. ∎
As with projections, we have a trivial, though still useful, bound for the complexity of distances.
Lemma 22.
Let $A\subseteq\mathbb{N}$, $x,y\in\mathbb{R}^{2}$, $e=\frac{x-y}{|x-y|}$, $\varepsilon>0$, and $t<r$. Suppose $r$ is sufficiently large. Then
$K^{A,x}_{r,r,t}(y\mid|x-y|,y)\leq r-t+\epsilon r$.
Proof.
Let
$C=\{z\in B_{2^{-t}}(y)\mid||x-z|-|x-y||<2^{-r}\}$.
Note that $C$ can be covered by $O(2^{r-t})$ dyadic cubes with side-length $2^{-r}$. Moreover, this covering is computable given $y$ to precision $t$ and given $x$ as an oracle, and the conclusion follows. ∎
We say that a partition $\mathcal{P}=\{[a_{i},a_{i+1}]\}_{i=0}^{k}$ of intervals with disjoint interiors is good if $[1,r]=\cup_{i}[a_{i},a_{i+1}]$ and it satisfies the following conditions.
  1. (1)
    $[a_{i},a_{i+1}]$ is either $\sigma$-yellow or $\sigma$-teal,
  2. (2)
    $a_{i+1}\leq 2a_{i}$, for every $i$ and
  3. (3)
    $a_{i+2}>2a_{i}$ for every $i<k$.
Essentially the same proof as Lemma 18 shows that a good partition gives a bound on the complexity of $K^{A,x}_{r}(y\mid|x-y|)$.
Lemma 23.
Let $\mathcal{P}$ be a good partition. Then
$K^{A,x}_{r}(y\mid|x-y|)\leq\epsilon r+\sum\limits_{i\in\mathbf{Bad}}\min\{K^{A}_{r-t}(y\mid y)-\sigma(r-t),r-t\}$,
where
Bad =$=\{i\leq k\mid[a_{i},a_{i+1}]\notin T_{\sigma}\}$.
4.1. Constructing a better partition
We now fix a $A\subseteq\mathbb{N}$, $\sigma\in(0,1)$, $x,y$ and $e:=\frac{y-x}{|x-y|}$ satisfying
  • (C1)
    $\dim^{A}(x),\dim^{A}(y)>\sigma$
  • (C2)
    $K^{A,x}_{r}(e)\geq\sigma r-O(\log r)$ for all $r$.
  • (C3)
    $K^{A,x}_{r}(y)\geq K^{A}_{r}(y)-O(\log r)$ for all sufficiently large $r$.
  • (C4)
    $K^{A}_{t,r}(e\mid y)\geq\sigma t-O(\log r)$ for all sufficiently large $r$ and $t\leq r$.
Let $D=\max\{\operatorname{Dim}^{A}(x),\operatorname{Dim}^{A}(y)\}$. Let $\alpha=\min\{D,1+\sigma\}$.
Constructing the partition $\mathcal{P}$: We define the sequence $\{r_{i}\}$ inductively. To begin, we set $r_{0}=r$. Having defined the sequence up to $r_{i}$, we choose $r_{i+1}$ as follows. Let $a\leq r_{i}$ be the minimal real such that $[a,r_{i}]$ can be written as the union of $\sigma$-yellow intervals whose lengths are at most doubling. If $a<r_{i}$, then we set $r_{i+1}=a$. In this case, we will refer to $[r_{i+1},r_{i}]$ as a yellow interval of $\mathcal{P}$.
Otherwise, let $t_{i}<r_{i}$ be the maximum of all reals $t<r$ such that
(17) $f(t)=f(r_{i})-\sigma(r_{i}-t)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t).$
Let $t_{i}^{\prime}<r_{i}$ be the largest real $t<r$ such that $f(r_{i})=f(t)+\sigma(r_{i}-t)$. Note that such a $t^{\prime}$ must exist. We then set
(18) $r_{i+1}=\max\{t_{i},t_{i}^{\prime}\}.$
Note that, in this case, $[r_{i+1},r_{i}]$ is $\sigma$-teal. We therefore refer these intervals as teal intervals of $\mathcal{P}$.
We begin by observing that our partition is well defined.
Observation 24.
Suppose that $r_{i}\leq r$ and $r_{i}>C$, for some fixed constant $C$ depending only on $y$. Then there is at least one $t$ such that
$f(t)=f(r_{i})-\sigma(r_{i}-t)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t).$
Proof.
We first note that $f(0)=O(1)$, and so, for $t=0$,
$f(r_{i})-\sigma(r_{i}-t)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t)$ $>f(r_{i})-\sigma r_{i}$
$>f(0),$
since $\dim^{A}(y)>\sigma$.
We also see that, at $t=r_{i}$,
$f(r_{i})-\sigma(r_{i}-t)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t)<f(r_{i}),$
and the conclusion follows. ∎
Having established that our partition is well-defined, we now show that the complexity of the distance has good growth on the yellow intervals.
Lemma 25.
Let $\varepsilon>0$. Suppose that $[r_{i+1},r_{i}]\in\mathcal{P}$ is a yellow interval. Then,
$K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)\geq\sigma(r_{i}-r_{i+1})-\varepsilon r_{i}.$
Proof.
By assumption, $[r_{i+1},r_{i}]$ is the union of of $\sigma$-yellow intervals $[a_{j+1},a_{j}]$ such that $a_{j}\leq 2a_{j+1}$. By a simple greedy strategy we can construct a partition $P_{1}=\{[b_{k+1},b_{k}]\}$ of $[r_{i+1},r_{i}]$ such that, for every $k$, $[b_{k+1},b_{k}]$ is yellow, $b_{k}\leq 2b_{k+1}$ and $b_{k+2}>2b_{k}$. That is, $P_{1}$ is a good partition of $[r_{i+1},r_{i}]$. The conclusion then follows from Lemma 23. ∎
We now turn to the growth rate of the complexity of $|x-y|$ on the teal intervals of our partition. This is somewhat involved, and will need several auxiliary results.
Lemma 26.
If $[r_{i+1},r_{i}]\in\mathcal{P}$ is teal, then $r_{i+1}\leq\frac{r_{i}}{2}$.
Proof.
Suppose that $[r_{i+1},r_{i}]\in\mathcal{P}$ is teal. Then, by the construction of $\mathcal{P}$, $[t,r_{i}]$ is not $\sigma$-yellow, for any $\frac{r_{i}}{2}\leq t<r_{i}$. This immediately implies that $t^{\prime}_{i}<\frac{r_{i}}{2}$. Moreover, for any $t>\frac{r_{i}}{2}$, we see that
$f(t)-f(r_{i})+\sigma(r_{i}-t)-\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t)$ $\geq\sigma(t-r_{i})+\sigma(r_{i}-t)-\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t)$
$=\frac{(\alpha-\sigma)\sigma}{\alpha}(2t-r_{i})$
$>0,$
implying that $t_{i}\leq\frac{r_{i}}{2}$, and the conclusion follows. ∎
We will need a lower bound on the left endpoint of teal intervals.
Lemma 27.
Let $\varepsilon>0$. Suppose that $r_{i+1}$ is sufficiently large and $[r_{i+1},r_{i}]\in\mathcal{P}$ is a teal interval. Then
$r_{i+1}\geq\frac{(\alpha-\sigma)\sigma}{\alpha D+\alpha\sigma-2\sigma^{2}}r_{i}-\varepsilon r_{i}$
Proof.
Assume that $r_{i+1}$ is large enough that, for all $s>r_{i+1}$,
$ds-\varepsilon^{\prime}s\leq f(s)\leq Ds+\varepsilon^{\prime}s$
for some $\varepsilon^{\prime}$ to be determined. If $K^{A}_{r_{i+1}}(y)=K^{A}_{r_{i}}(y)-\sigma(r_{i}-r_{i+1})$, then the conclusion follows by the definition of our partition. Otherwise, we have
(19) $K^{A}_{r_{i+1}}(y)\geq K^{A}_{r_{i}}(y)-\sigma(r_{i}-r_{i+1})+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2r_{i+1})-\varepsilon^{\prime}r_{i}.$
Therefore,
$Dr_{i+1}$ $\geq D_{y}r_{i+1}$
$\geq K^{A}_{r_{i+1}}(y)-\varepsilon^{\prime}r_{i+1}$
$\geq K^{A}_{r_{i}}(y)-\sigma(r_{i}-r_{i+1})+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2r_{i+1})-\varepsilon^{\prime}r_{i}$
$\geq\sigma r_{i+1}+\frac{(\alpha-\sigma)\sigma}{D}(r_{i}-2r_{i+1})-2\varepsilon^{\prime}r_{i}$
$\geq\frac{(\alpha-\sigma)\sigma}{\alpha}r_{i}+\frac{2\sigma^{2}-\sigma\alpha}{\alpha}r_{i+1}-2\varepsilon^{\prime}r_{i}$
Rearranging, choosing $\varepsilon^{\prime}$ appropriately, and simplifying yields
(20) $r_{i+1}\geq\frac{(\alpha-\sigma)\sigma}{\alpha D+\alpha\sigma-2\sigma^{2}}r_{i}-\varepsilon r_{i}.$
With the previous lemmas, we can prove a strong lower bound on the slope of the complexity of $y$ on teal intervals.
Lemma 28.
Let $\varepsilon>0$ be given and suppose that $[r_{i+1},r_{i}]\in\mathcal{P}$ is a teal interval and $r_{i}$ is sufficiently large. Then
(21) $\frac{K^{A}_{r_{i},r_{i+1}}(y\mid y)}{r_{i}-r_{i+1}}\geq\frac{\sigma^{2}(D+\alpha-2\sigma)}{\alpha D-\sigma^{2}}-\varepsilon$
Proof.
Recall that we chose $r_{i+1}$ to be
$r_{i+1}=\max\{t,t^{\prime}\}$,
where $t^{\prime}$ is the largest real such that $[t^{\prime},r_{i}]$ is green, and $t$ is the largest real such that
$f(t)=f(r_{i})-\sigma(r_{i}-t)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2t).$
If $r_{i+1}=t^{\prime}$, then
$K^{A}_{r_{i+1}}(y)=K^{A}_{r_{i}}(y)-\sigma(r_{i}-r_{i+1}),$
and the conclusion holds trivially.
We now assume that $r_{i+1}=t$. Then we calculate
$K^{A}_{r_{i}}(y)-K^{A}_{r_{i+1}}$ $=\sigma(r_{i}-r_{i+1})-\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2r_{i+1})$
$=\frac{\sigma^{2}}{\alpha}r_{i}-\frac{2\sigma^{2}-\alpha\sigma}{\alpha}r_{i+1}.$
Hence, it suffices to show that
(22) $\frac{\sigma^{2}}{\alpha}r_{i}-\frac{2\sigma^{2}-\alpha\sigma}{\alpha}r_{i+1}\geq\left(\frac{\sigma^{2}(D+\alpha-2\sigma)}{\alpha D-\sigma^{2}}-\varepsilon\right)(r_{i}-r_{i+1})$
Routine calculation shows that this is true, since
(23) $r_{i+1}\geq\frac{(\alpha-\sigma)\sigma}{\alpha D+\alpha\sigma-2\sigma^{2}}r_{i}-\varepsilon r_{i}.$
by Lemma 27.
We are now able to give strong lower bounds on the complexity $K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)$ for any teal interval $[r_{i+1},r_{i}]\in\mathcal{P}$. To do so, we will use the following from [27].
Lemma 29.
Let $x,y\in\mathbb{R}^{2}$ and $r\in\mathbb{N}$. Let $z\in\mathbb{R}^{2}$ such that $|x-y|=|x-z|$. Then for every $A\subseteq\mathbb{N}$,
(24) $K^{A}_{r}(z)\geq K^{A}_{t}(y)+K^{A}_{r-t,r}(x\mid y)-K_{r-t}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r),$
where $e^{\prime}=\frac{y-z}{|y-z|}$ and $t=-\log|y-z|$.
Lemma 30.
Let $\varepsilon>0$. Suppose that $[r_{i+1},r_{i}]\in\mathcal{P}$ is a teal interval and $r_{i+1}$ is sufficiently large. Then,
(25) $K^{A,x}_{r_{i},r_{i},r_{i+1}}(y\mid|x-y|,y)\leq\varepsilon r_{i}.$
In particular, $K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)\geq K^{A,x}_{r_{i},r_{i+1}}(y\mid y)-\varepsilon r_{i}$.
Proof.
Let $\varepsilon\in\mathbb{Q}$ be sufficiently small, assume $r_{i+1}$ is sufficiently large. Let $\eta$ be the rational such that $\eta r_{i}=K^{A}_{r_{i}}(y)-4\varepsilon r_{i}$. Let $G=D^{A}(y,r,\eta)$ be the oracle of Lemma 9. Then, relative to the oracle $(A,G)$, the first condition of Lemma 20 holds. So, we need to show that for $z\in B_{2^{-r_{i+1}}}(y)$ such that $|x-y|=|x-z|$,
$K^{A,G}_{r}(z)\geq\eta r_{i}+\min\{\varepsilon r_{i},\sigma(r_{i}-s)-\varepsilon r_{i}\}$
where $s=-\log|y-z|.$
As observed in [27], if $s\geq\frac{r_{i}}{2}-\log r_{i}$, then $|p_{e}y-p_{e}z|\leq{r_{i}}^{2}2^{-r_{i}}$. Hence, applying Lemma 12 and using the properties of our direction $e$,
$K^{A,G}_{r_{i}}(z)\geq K^{A,G}_{s}(y)+\sigma(r_{i}-s)-\frac{\varepsilon}{2}r_{i}-O(\log r_{i}).$
Using the properties of $G$ and the fact that $[r_{i+1},r_{i}]$ is teal establishes the desired bound in this case, so we assume $s\leq\frac{r_{i}}{2}-\log r_{i}$. First, apply Lemma 29 to obtain
$K^{A,G}_{r}(z)\geq K^{A,G}_{t}(y)+K^{A,G}_{r-t,r}(x\mid y)-K^{A,G}_{r-t}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
We want to apply Theorem 11, so we check that its conditions are satisfied with respect to x, $e^{\prime},\varepsilon,C,t=s,$ and $r=r_{i}-s$. The first condition of the theorem is satisfied using (C1), and the third is satisfied using (C2), so it remains to verify the bound on $t$. We know $r_{i+1}<s$, and can additionally apply Lemma 27 to obtain
$s$ $\geq\dfrac{(\alpha-\sigma)\sigma}{\alpha D+\alpha\sigma-2\sigma^{2}}r_{i}-\varepsilon r_{i}$
$\geq\dfrac{(\alpha-\sigma)\sigma}{\alpha D-\sigma^{2}}(r_{i}-s)-\varepsilon r_{i}$
$\geq\dfrac{\sigma(\sigma+1-\alpha)}{\alpha+\sigma+\sigma^{2}-\alpha\sigma}(r_{i}-s)$
where the last inequality follows if $\varepsilon$ is small enough depending on $\sigma$ and $D$, which is no significant encumbrance. So, we may apply Theorem 11, the properties of $G$, and (C3) with symmetry of information to obtain
$K^{A,G}_{r_{i}}(z)$ $\geq K^{A,G}_{s}(y)+K^{A,G}_{r_{i}-s,r_{i}}(x\mid y)-K^{A,G}_{r_{i}-s}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
$\geq K^{A}_{s}(y)+K^{A}_{r_{i}-s}(x)-K^{A}_{r_{i}-s}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
$\geq K^{A}_{s}(y)+K^{A}_{r_{i}-s}(x)-\left(K^{A}_{r_{i}-s}(x)-\sigma(r_{i}-s)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2s)\right)-\varepsilon r_{i}-O(\log r)$
$\geq K^{A}_{s}(y)+\sigma(r_{i}-s)-\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2s)-\varepsilon r_{i}-O(\log r)$
We may use (19) and the teal property to get the bound
$K^{A}_{s}(y)\geq K^{A}_{r_{i}}(y)-\sigma(r_{i}-s)+\frac{(\alpha-\sigma)\sigma}{\alpha}(r_{i}-2s)-\varepsilon r_{i}$
which in conjunction with the previous equation gives
$K^{A,G}_{r_{i}}(z)\geq K_{r_{i}}^{A}(y)-3\varepsilon r_{i}\geq\eta r_{i}+\varepsilon r_{i}$
Thus the second condition of Lemma 20 is satisfied, and applying it gives the desired bound. ∎
4.2. Main effective theorem
We are now in a position to prove our main effective theorem. This is the effective analog of Theorem 2, the main theorem of this paper.
Theorem 31.
Suppose that $x,y\in\mathbb{R}^{2}$, $e=\frac{y-x}{|y-x|}$, $10<\sigma<$ 1 and $A\subseteq\mathbb{N}$ satisfy the following.
  • (C1)
    $\dim^{A}(x),\dim^{A}(y)>\sigma$
  • (C2)
    $K^{A,x}_{r}(e)>\sigma r-O(\log r)$ for all $r$.
  • (C3)
    $K^{A,x}_{r}(y)\geq K^{A}_{r}(y)-O(\log r)$ for all sufficiently large $r$.
  • (C4)
    $K^{A}_{t,r}(e\mid y)>\sigma t-O(\log r)$ for all sufficiently large $r$ and $t\leq r$.
Then
$\dim^{x,A}(|x-y|)\geq\sigma\left(1-\frac{\alpha D-\sigma(D+\alpha-\sigma)}{(\sigma+1)(\alpha D-\sigma^{2})-\sigma^{2}(\alpha+D-2\sigma)}\right),$
where $D=\max\{D_{x},D_{y}\}$ and $\alpha=\min\{D,1+\sigma\}$.
Proof.
Let $\sigma^{\prime}\in\mathbb{Q}$ satisfying $0<\sigma^{\prime}<\sigma$ and let $\varepsilon>0$. Let $r$ be sufficiently large. Let $\mathcal{P}=\{[r_{i+1},r_{i}]\}$ be the partition of $[1,r]$ defined in the previous section. Let $L$ be the total length of the $\sigma^{\prime}$-yellow intervals. Recall that if $[r_{i+1},r_{i}]$ is yellow, then we have that
$K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)\geq\sigma^{\prime}(r_{i}-r_{i+1})-\varepsilon r_{i}$
By applications of the symmetry of information and Lemmas 25 and 30 , for sufficiently large $r$
$K^{A,x}_{r}(|x-y|)$ $=\sum\limits_{i\in Y}K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)+\sum\limits_{i\in Y^{C}}K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)$
$\geq\sigma^{\prime}L-\frac{\varepsilon}{3}r+\sum\limits_{i\in Y^{C}}K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)$
$\geq\sigma^{\prime}L-\frac{2\varepsilon}{3}r+\sum\limits_{i\in Y^{C}}K^{A,x}_{r_{i},r_{i+1}}(y\mid y)$
$\geq\sigma^{\prime}L+\frac{\sigma^{2}(D+\alpha-2\sigma)}{\alpha D-(\sigma^{\prime})^{2}}(r-L)-\varepsilon r.$
We also have the bound
$d_{y}r$ $\leq K^{A,x}_{r}(y)+\frac{\varepsilon}{3}$
$=\sum\limits_{i\in Y}K^{A,x}_{r_{i},r_{i+1}}(y\mid y)+\sum\limits_{i\in Y^{C}}K^{A,x}_{r_{i},r_{i+1}}(y\mid y)$
$\leq L+\sum\limits_{i\in Y}K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)+\frac{2\varepsilon}{3}$
$\;\;\;\;\;\;\;\;\;\;\;+\sum\limits_{i\in Y^{C}}K^{A,x}_{r_{i},r_{i+1}}(|x-y|\mid|x-y|)$
$\leq L+K^{A,x}_{r}(|x-y|)+\varepsilon r.$
Hence,
(26) $K^{A,x}_{r}(|x-y|)\geq\max\{\sigma^{\prime}L+\frac{(\sigma^{\prime})^{2}(D+\alpha-2\sigma^{\prime})}{\alpha D-(\sigma^{\prime})^{2}}(r-L),\sigma^{\prime}r-L\}-\varepsilon r.$
The first term is increasing in $L$, and the second term is decreasing in $L$, so we can set them equal to find the minimum over all $L$, which yields
(27) $K^{x,A}_{r}(|x-y|)\geq\sigma^{\prime}\left(1-\frac{\alpha D-\sigma^{\prime}(D+\alpha-\sigma^{\prime})}{(\sigma^{\prime}+1)(\alpha D-(\sigma^{\prime})^{2})-(\sigma^{\prime})^{2}(\alpha+D-2\sigma^{\prime})}\right)r-\varepsilon r.$
Since we can take $\sigma^{\prime}$ arbitrarily close to $\sigma$ and $\varepsilon$ as small as desired, the conclusion follows.
5. Reduction to the classical bounds
Recall the conditions on points in our main effective theorem. For an oracle $A\subseteq\mathbb{N}$ we want $x,y\in\mathbb{R}^{2}$ and $e=\frac{y-x}{|x-y|}$ to satisfy.
  • (C1)
    $\dim^{A}(x),\dim^{A}(y)>\sigma$
  • (C2)
    $K^{A,x}_{r}(e)>\sigma r-O(\log r)$ for all $r$.
  • (C3)
    $K^{A,x}_{r}(y)\geq K^{A}_{r}(y)-O(\log r)$ for all sufficiently large $r$.
  • (C4)
    $K^{A}_{t,r}(e\mid y)>\sigma t-O(\log r)$ for all sufficiently large $r$ and $t\leq r$.
We need to prove the existence of points satisfying these conditions.
Let $K,t\geq 0$, $c\in(0,1]$. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ with $\text{spt}(\mu)=:X$ and $\text{spt}(\nu)=:Y$. We say that $(\mu,\nu)$ has $(t,K,c)$-thin tubes if there is a Borel set $G\subset X\times Y$ with $(\mu\times\nu)(G)\geq c$ with the following property. If $x\in X$, then
$\nu(T\cap G|_{x})\leq Kr^{t}$ for all $r>0$ and all $r$-tubes $T$ containing $x$.
We say that $(\mu,\nu)$ has $t$-thin tubes if $(\mu,\nu)$ has $(t,K,c)$-thin tubes for some $K,c>0$.
We will need the following corollary proved by Orponen, Shmerkin and Wang [22] (Corollary 2.21).
Corollary 32.
Let $s\in(0,1]$ and let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{2})$ be measures which satisfy the $s$-dimensional Frostman condition $\mu(B(x,r))\lesssim r^{s}$, $\nu(B(x,r))\lesssim r^{s}$, for all $x\in\mathbb{R}^{2}$ and $r>0$. Assume $\mu(\ell)\nu(\ell)<1$ for every line $\ell$. Then $(\mu,\nu)$ has $\sigma$-thin tubes for all $0\leq\sigma<s$.
We use the existence of thin tubes to establish the existence of points $x,y$ satisfying properties (C1)-(C3)
Lemma 33.
Let $A\subseteq\mathbb{N}$, $X,Y\subseteq\mathbb{R}^{2}$ be compact sets, $s\in(0,1]$ and $\mu,\nu\in\mathcal{P}(\mathbb{R}^{2})$ be measures whose support is $X$ and $Y$, respectively, which satisfy the $s$-dimensional Frostman condition. Assume $\mu(\ell)\nu(\ell)<1$ for every line $\ell$ and $\nu$ is computable with respect to $A$. Then, for every $\sigma<\min\{s,1\}$,
$\dim_{H}\{x\in X:\exists y\in Y\text{ satisfying (C1)-(C3)}\}\geq\sigma$
Proof.
By Corollary 32, $(\mu,\nu)$ has $\sigma^{\prime}$-thin tubes, for every $\sigma<\sigma^{\prime}<s$. Let $G\subseteq X\times Y$ be the Borel set in the definition of $\sigma^{\prime}$-thin tubes. Since $(\mu\times\nu)(G)\geq c$, the set
$X_{1}=\{x\in X\mid\nu(G|_{x})>0\}$
has positive $\mu$-measure. The Hausdorff dimension of $X_{1}$ is at least $s>\sigma$, so, by the point to set principle, the set
$X_{2}=\{x\in X_{1}\mid\dim^{A}(x)>\sigma\}$
has Hausdorff dimension $s$.
Let $x$ be any point in $X_{2}$, and let $G^{\prime}\subseteq G|_{x}$ be the set of points in $G|_{x}$ satisfying (C1) and (C3). Then, since $\nu(G|_{x})>0$, by [27] (Lemma 36), $\nu(G^{\prime})=\nu(G|_{x})>0$.
Since $(\mu,\nu)$ has $\sigma^{\prime}$-thin tubes, we must have
$\nu(T\cap G^{\prime})\leq Kr^{\sigma^{\prime}}$ for all $r>0$ and all $r$-tubes $T$ containing $x$.
In particular, $\dim_{H}(\pi_{x}G^{\prime})\geq\sigma^{\prime}$. By the point-to-set principle, this implies that there is a direction $e\in\pi_{x}G^{\prime}$ such that $\dim^{A,x}(e)>\sigma$. Let $e$ be such a direction, and let $y\in G^{\prime}$ such that $e=\frac{y-x}{|x-y|}$. Then this choice of $x$ and $y$ satisfies (C1)-(C3). ∎
Now, we prove that (C1)-(C3) imply (C4)
Lemma 34.
Assume $A\subseteq\mathbb{N}$, $x,y\in\mathbb{R}^{2}$, $e=\frac{y-x}{|x-y|}$ and $\sigma\in(0,1)$ satisfy (C1)-(C3). Then $x$ and $y$ also satisfy (C4), i.e.,
$K^{A}_{t,r}(e\mid y)>\sigma t-O(\log r)$,
for all sufficiently large $r\in\mathbb{N}$ and $t\leq r$.
Proof.
Let $\varepsilon>0$, $r\in\mathbb{N}$ be sufficiently large and $t\leq r$. By the symmetry of information, and the fact that $e=\frac{y-x}{|x-y|}$,
$K^{A}_{t,r}(x\mid y)$ $\leq K^{A}_{t,r}(e\mid y)+K^{A}_{t,t,r}(x\mid e,y)+O(\log r)$
$\leq K^{A}_{t,r}(e\mid y)+K^{A}_{t}(x\mid p_{e^{\bot}}x,e^{\bot})+O(\log r).$
We note that $e^{\bot}$ and $e$ are Lipschitz-bi-computable, and so $K_{r}(e)=K_{r}(e^{\bot})+O(\log r)$. By a pointwise analog of Marstrand’s projection theorem, [^3] Essentially proven in [17], but also a result of the projection bound in section 3 arising from a one interval partition
$K^{A}_{t}(x\mid p_{e^{\bot}}x,e^{\bot})<K^{A}_{t}(x)-\sigma t+O(\log t),$
and so
(28) $K^{A}_{t,r}(e\mid y)\geq\sigma t+K^{A}_{t,r}(x\mid y)-K^{A}_{t}(x)-O(\log r).$
By condition (C3), and the symmetry of information again,
$K^{A}_{t,r}(x\mid y)$ $=K^{A}_{t}(x)+K^{A}_{r,t}(y\mid x)-K^{A}_{r}(y)-O(\log r)$
$\geq K^{A}_{t}(x)+K^{A,x}_{r}(y)-K^{A}_{r}(y)-O(\log r)$
$\geq K^{A}_{t}(x)-O(\log r),$
and the conclusion follows. ∎
We are now in a position to prove our classical bounds.
Theorem 35.
Suppose $X$ and $Y$ are analytic sets such that $\dim_{H}(X),\dim_{H}(Y)>\sigma$, and $\dim_{P}(X),\dim_{P}(Y)\leq D$. Then for a set of $x\in X$ of Hausdorff dimension at least $\sigma$,
$\dim_{H}(\Delta_{x}(Y))\geq\sigma\left(1-\frac{\alpha D-\sigma(D+\alpha-\sigma)}{(\sigma+1)(\alpha D-\sigma^{2})-\sigma^{2}(\alpha+D-2\sigma)}\right)$
Where $\alpha=\min\{1+\sigma,D\}$
Proof.
Through standard results in geometric measure theory (see, e.g., [19]), there are compact subsets $X_{1}\subseteq X$ and $Y_{1}\subseteq Y$ such that $\dim_{H}(X_{1}),\dim_{H}(Y_{1})>\sigma$. To simplify notation, we denote these compact subsets $X$ and $Y$, respectively. In addition, there are measures $\mu,\nu\in\mathcal{P}(\mathbb{R}^{2})$ whose support is $X$ and $Y$, and which satisfy the $s$-dimensional Frostman condition, for some $s>\sigma$.
If, for some line $\ell$, $\mu(\ell)\nu(\ell)=1$, then the conclusion holds trivially. So we assume that, for every line $\ell$, $\mu(\ell)\nu(\ell)<1$.
Let $A$ the join of the following oracles: an oracle relative to which $Y$ is effectively compact, a packing oracle for $X$, a packing oracle for $Y$, and an oracle relative to which $\nu$ is computable. Applying Lemma 33 gives the existence of some $x$ and $y$ satisfying conditions (C1)-(C3) relative to $A$. By Lemma 34, this pair also satisfies condition (C4). By the properties of $A$, $\operatorname{Dim}^{A}(x),\operatorname{Dim}^{A}(y)\leq D$, hence we can apply Theorem 31 and conclude that
$\dim^{A,x}(|x-y|)\geq\sigma\left(1-\frac{\alpha D-\sigma(D+\alpha-\sigma)}{(\sigma+1)(\alpha D-\sigma^{2})-\sigma^{2}(\alpha+D-2\sigma)}\right)$
We complete the proof using two observations in [27]
  • If $Y$ is effectively compact relative to $A$, $A$ is a Hausdorff oracle for $Y$.
  • If $Y$ is effectively compact relative to $A$, then $\Delta_{x}Y$ is effectively compact relative to $(A,x)$.
By our choice of $A$, then, $(A,x)$ is a Hausdorff oracle for $\Delta_{x}Y$, and so by the point to set principle
(29) $\dim_{H}(\Delta_{x}Y)\geq\sigma\left(1-\frac{\alpha D-\sigma(D+\alpha-\sigma)}{(\sigma+1)(\alpha D-\sigma^{2})-\sigma^{2}(\alpha+D-2\sigma)}\right).$
Corollary 36.
Suppose $X$ and $Y$ are analytic sets such that $\dim_{H}(X),\dim_{H}(Y)>\sigma$. Then for a set of $x\in X$ of Hausdorff dimension at least $\sigma$,
$\dim_{H}(\Delta_{x}Y)\geq\sigma\left(1-\frac{2-\sigma}{2+4\sigma-2\sigma^{2}}\right)$
Proof.
It is easy to verify that the function in the bound (29) is deceasing in $D$ for all $\sigma\in[0,1]$. Hence, we may assume $D=2$, which implies $\alpha=1+\sigma$, and the conclusion follows. ∎
Theorem 2.
Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set, $d:=\dim_{H}(E)\leq 1$ and $D:=\dim_{P}(E)$. Then,
$\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{\alpha D-d(D+\alpha-d)}{(d+1)(\alpha D-d^{2})-d^{2}(\alpha+D-2d)}\right)$,
where $\alpha=\min\{1+d,D\}$. In particular,
$\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{2-d}{2+4d-2d^{2}}\right)$
Proof.
Since $E$ is analytic, we can find disjoint, analytic subsets $X,Y\subseteq E$ such that $\dim_{H}(X)=\dim_{H}(Y)=d$. By Theorem 35, for every $\sigma<d$, there is a point $x\in E$ such that
(30) $\dim_{H}(\Delta_{x}Y)\geq\sigma\left(1-\frac{\alpha D-\sigma(D+\alpha-\sigma)}{(\sigma+1)(\alpha D-\sigma^{2})-\sigma^{2}(\alpha+D-2\sigma)}\right),$
and the conclusion follows. ∎
6. Universal pin sets
Let $\mathcal{C}$ be a class of subsets of $\mathbb{R}^{2}$. Recall that $X$ is weakly universal for pinned distances for $\mathcal{C}$ if, for every $Y\in\mathcal{C}$
$\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))=\min\{\dim_{H}(Y),1\}.$
and universal for pinned distances for $\mathcal{C}$ if for every $Y\in\mathcal{C}$ there exists some $x\in X$ such that
$\dim_{H}(\Delta_{x}(Y))=\min\{\dim_{H}(Y),1\}.$
In [8], the authors established the following result.
Theorem 37.
Let $X\subseteq\mathbb{R}^{2}$ be such that $1<d_{x}<\dim_{H}(X)=\dim_{P}(X)$ and let $Y\subseteq\mathbb{R}^{2}$ be analytic and satisfy $1<d_{y}<\dim_{H}(Y)$. Then there is some $F\subseteq X$ such that,
$\dim_{H}(\Delta_{x}Y)=1$
for all $x\in F$. Moreover, $\dim_{H}(X\setminus F)<\dim_{H}(X)$.
In the language of universal sets, this theorem immediately implies that every weakly regular set with Hausdorff dimension greater than one is universal for the class of analytic sets in the plane having dimension more than 1. The regularity of the pin set was essential, as it ensured that for typical pins, there was a nearly optimal projection bound. We will prove two results concerning the case where $Y$ has dimension no more than 1; combined with the above these will give the existence of (weakly) universal sets for classes with no dimension restriction.
First, we will consider the case that $X$ is weakly regular. In this case, we can guarantee that typical points are arbitrarily close to (effectively) regular, but not necessarily regular (in the sense of having the same effective dimension relative to an appropriate oracle). As Theorem 37 indicates, if $X$ and $Y$ both have dimension more than 1, this still is enough for the full strength conclusion of universality. However, if $Y$ has dimension at most $1$, we can only achieve weak universality. This is the content of the first subsection.
In the second subsection, we assume additional regularity in $X$, namely that it is AD regular [^4] Full Ahlfors-David regularity isn’t necessary, it is possible to weaken the assumption slightly. However, we will require AD regularity to simplify the exposition.. In this case, we are guaranteed the existence of points that are extremely regular, with their complexity functions only deviating from $\dim_{H}(x)r$ by a logarithmic term. This enables us to show an optimal effective theorem. We only consider the case that $\dim_{H}(X)>1$ so we can employ a somewhat different reduction, though one that still depends on [22].
6.1. Weakly regular universal sets
We begin with a projection bound for the weakly regular case.
Proposition 38.
Let $A\subseteq\mathbb{N}$, $x\in\mathbb{R}^{2}$, $e\in\mathcal{S}^{1}$, $\sigma\in\mathbb{Q}\cap(0,1)$, $\varepsilon\in\mathbb{Q}^{+}$, $C,C^{\prime}\in\mathbb{N}$, and $t,r\in\mathbb{N}$. Suppose that $r$ is sufficiently large, and that the following hold.
  1. (P1)
    $\sigma<\dim^{A}(x)=:d_{x}$.
  2. (P2)
    $t\geq\frac{r}{C}$.
  3. (P3)
    $K^{x,A}_{s}(e)\geq\sigma s-C^{\prime}\log s$, for all $s\leq t$.
Then there exists some $\varepsilon_{x}$ depending only on $d_{x}$ and $C$ such that $\operatorname{Dim}^{A}(x)-\dim^{A}(x)<\varepsilon_{x}$ implies that
$K^{A}_{r}(x\,|\,p_{e}x,e)\leq K^{A}_{r}(x)-\sigma r+\varepsilon r.$
Proof.
Let $C$ and $\varepsilon$ be given. Assume $r$ is large enough that if $\sqrt{r}<s$
$d_{x}s-\varepsilon^{\prime}s\leq K_{s}^{A}(x)+d_{x}s-\varepsilon^{\prime}s$
where $\varepsilon^{\prime}>\varepsilon_{x}$ and may further depend on $C$, $d_{x}$, and $\sigma$. Let $t>\frac{r}{C}$ be given and let $\mathcal{P}$ be the partition of $[1,r]$ into red, $\sigma$-green, and blue intervals of Lemma 19. We will show that after a small precision, all the intervals in this partition are green or red, which will allow us to apply Lemma 18 to a partition that almost entirely consists of yellow intervals.
By the properties of $\mathcal{P}$, and the fact that $d_{x}>\sigma$, this follows if there are no red $\sigma$-green blue sequences in $\mathcal{P}$ after some small precision.
Any block of $\sigma$-green intervals in $\mathcal{P}$ always has length at least $t$, and the growth rate is exactly $\sigma$. Suppose $s\in[\sqrt{r},r]$ and $[s,s+u]$ is a green interval or union of adjacent green intervals. Then
$K_{s}^{A}(x)\leq d_{x}s+\varepsilon^{\prime}s$
while
$K_{s+u}^{A}(x)\geq d_{x}(s+u)-\varepsilon^{\prime}(s+u)$
By our assumption, $K$ grows by $\sigma u$ on $[s,s+u]$, hence
$\sigma u$ $\geq(d_{x}-\varepsilon^{\prime})(s+u)-(d_{x}+\varepsilon^{\prime})s$
$=d_{x}u-2\varepsilon^{\prime}s-\varepsilon^{\prime}u$
$\geq d_{x}u-2\varepsilon^{\prime}r-\varepsilon^{\prime}u.$
Using the fact that $t\geq\frac{r}{C}$, this implies
$(d_{x}-\sigma-\varepsilon^{\prime})u\leq 2\varepsilon^{\prime}Ct$
So $u<t$ if
$\dfrac{2\varepsilon^{\prime}C}{d_{x}-\sigma-\varepsilon^{\prime}}<1.$
We are restricted to choosing $\varepsilon^{\prime}>\varepsilon_{x}$, but $\varepsilon_{x}$ can be chosen depending on $\sigma$, $d_{x}$, and $C$, so we may assume $\varepsilon^{\prime}$ is small enough that, for sufficiently large $r$, $u<t$ and consequently there are no red-green-blue sequences in $\mathcal{P}$.
The argument that, for $r$ sufficiently large, $[\sqrt{r},\frac{\varepsilon}{2}r]$ intersects a red interval proceeds in the same way as the proof of Proposition 40 in [8]. Hence, we omit it and observe the consequence that, for sufficiently large $r$, $[\frac{ve}{2}r,r]$ is covered entirely by red and green intervals. Using a simple greedy strategy, we can cover $[\frac{ve}{2}r,r]$ by no more than $2C$ yellow intervals. By Lemma 18 with respect to $\frac{\varepsilon}{2}$,
$K^{A}_{r}(x\mid p_{e}x,e)$ $\leq\frac{\varepsilon r}{2}+\sum\limits_{i\in\textbf{Bad}}\min\{K^{A}_{a_{i+1},a_{i}}(x\mid x)-\sigma(a_{i+1}-a_{i}),a_{i+1}-a_{i}\}$
$\leq\min\{K_{r}^{A}(x)-\sigma r,r\}+\varepsilon r$
Theorem 39.
Suppose that $x,y\in\mathbb{R}^{2}$, $e=\frac{y-x}{|y-x|}$, $\sigma\in\mathbb{Q}\cap(0,1)$, $\sigma^{\prime}>\sigma$, $\varepsilon>0$, and $A\subseteq\mathbb{N}$ satisfy the following.
  • (C1)
    $\dim^{A}(x)>\sigma^{\prime},\dim^{A}(y)>\sigma$
  • (C2)
    $K^{A,x}_{r}(e)>\sigma r-O(\log r)$ for all $r$.
  • (C3)
    $K^{A,x}_{r}(y)\geq K^{A}_{r}(y)-O(\log r)$ for all sufficiently large $r$.
  • (C4)
    $K^{A}_{t,r}(e\mid y)>\sigma t-O(\log r)$ for all sufficiently large $r$ and $t\leq r$.
Then there is a constant $\delta(\varepsilon,\sigma^{\prime})>0$, depending only on $\varepsilon$ and $s$ such that, if $\operatorname{Dim}^{A}(x)-\dim^{A}(x)<\delta(\varepsilon,\sigma^{\prime})$,
$\dim^{x,A}(|x-y|)\geq\sigma-\frac{10\varepsilon}{\sigma}.$
Proof.
Let $C=\frac{1}{10\varepsilon}$. Let $\delta(\varepsilon,\sigma^{\prime})$ be the real $\varepsilon_{x}$ in the statement of Proposition 38, and suppose that $\operatorname{Dim}^{A}(x)-\dim^{A}(x)<\delta(\varepsilon,\sigma^{\prime})$.
Let $\eta$ be a rational such that $\eta\leq\sigma-3\varepsilon$. Let $r$ be sufficiently large and let $t=\frac{r}{C}$. Let $G=D^{A}(r,y,\eta)$ be the oracle of Lemma 9. Then, relative to the oracle $(A,G)$, the first condition of Lemma 20 holds. So, we need to show that for $z\in B_{2^{-t}}(y)$ such that $|x-y|=|x-z|$,
$K^{A,G}_{r}(z)\geq\eta r+\min\{\varepsilon r,\sigma(r-s)-\varepsilon r\}$
where $s=-\log|y-z|.$
If $s\geq\frac{r}{2}-\log r$, then $|p_{e}y-p_{e}z|\leq{r}^{2}2^{-r}$. Lemma 12 shows that
$K^{A,G}_{r}(z)\geq K^{A,G}_{s}(y)+\sigma(r-s)-\frac{\varepsilon}{2}r-O(\log r).$
Using the properties of $G$ establishes the desired bound in this case, so we assume $s\leq\frac{r}{2}-\log r$. By Lemma 29 we see that
$K^{A,G}_{r}(z)\geq K^{A,G}_{s}(y)+K^{A,G}_{r-s,r}(x\mid y)-K^{A,G}_{r-s}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
It is straightforward to verify that the conditions of Proposition 38 are satisfied with respect to $x$, $e^{\prime},\varepsilon,C,t=s,$ and $r=r-s$. Applying Proposition 38 yields
$K^{A,G}_{r}(z)$ $\geq K^{A,G}_{s}(y)+K^{A,G}_{r-s,r}(x\mid y)-K^{A,G}_{r-s}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
$\geq K^{A}_{s}(y)+K^{A}_{r-s}(x)-K^{A}_{r-s}(x\mid p_{e^{\prime}}x,e^{\prime})-O(\log r)$
$\geq K^{A}_{s}(y)+K^{A}_{r-s}(x)-\left(K^{A}_{r-s}(x)-\sigma(r-s)+\varepsilon r\right)-\varepsilon r-O(\log r)$
$\geq K^{A}_{s}(y)+\sigma(r-s)-2\varepsilon r-O(\log r)$
$>\eta r-2\varepsilon r-O(\log r)$
establishing the second condition of Lemma 20. Therefore,
(31) $K^{A,x}_{r,t}(|x-y|\mid y)\geq\eta r-\frac{3\varepsilon}{\sigma}r-O(\log r).$
Hence, by definition of $\eta$,
(32) $\dim^{A,x}(|x-y|)\geq\sigma-\frac{10\varepsilon}{\sigma}.$
Theorem 40.
Let $X\subseteq\mathbb{R}^{2}$ be Borel and weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$. Then $X$ is universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$ satisfying $\dim_{H}(Y)<\dim_{H}(X)$.
Proof.
Let $\sigma<\dim_{H}(Y)<\dim_{H}(X)$, and let $\varepsilon>0$ be sufficiently small. Let $s$ be any real such that $\dim_{H}(Y)<s<\dim_{H}(X)$. Let $\delta=\delta(\varepsilon,s)$ be the real of Theorem 39. Let $X_{1}\subseteq X$ and $Y_{1}\subseteq Y$ be compact subsets such that
$\dim_{H}(X_{1})>\max\{\dim_{P}(X)-\delta,s\}$
and $\dim_{H}(Y_{1})>\sigma$. To simplify notation, we denote these compact subsets $X$ and $Y$, respectively. In addition, let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{2})$ whose support is $X$ and $Y$. We assume $\mu$ and $\nu$ satisfy the $s_{X}$-dimensional and $s_{Y}$-dimiensional, respectively, Frostman condition where $s_{X}>\max\{\dim_{P}(X)-\delta,s\}$ and $s_{Y}>\sigma$.
Let $A$ the join of the following oracles: an oracle relative to which $Y$ is effectively compact, a packing oracle for $X$, and an oracle relative to which $\nu$ is computable. The proof of Lemma 33 gives the existence of some $x$ and $y$ satisfying conditions (C1)-(C3) relative to $A$. Lemma 34 proves $x$ and $y$ also satisfy condition (C4). By Theorem 39
$\dim^{A,x}(|x-y|)\geq\sigma-\frac{10\varepsilon}{\sigma}$
Since $Y$ is effectively compact relative to $A$, we see that
$\dim_{H}(\Delta_{x}(Y))\geq\sigma-\frac{10\varepsilon}{\sigma}$.
Since $\varepsilon$ can be arbitrarily small, we have
$\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))\geq\sigma$.
Finally, since $\sigma$ can be arbitrarily close to $\dim_{H}(Y)$, the conclusion follows. ∎
Combining this result with Theorem 37 immediately gives the follows corollary.
Corollary 41.
Let $X\subseteq\mathbb{R}^{2}$ be Borel and weakly regular with $\dim_{H}(E)=\dim_{P}(E)>1$. Then $X$ is weakly universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$.
6.2. AD-regular universal sets
We begin by defining AD-regular sets and proving several pointwise results for them. Let $E\subseteq\mathbb{R}^{n}$ be a compact set, $C\geq 1$ and $\alpha\geq 0$. We say that $E$ is $(\alpha,C)$-AD regular if
(33) $C^{-1}r^{\alpha}\leq\mathcal{H}^{\alpha}(E\cap B(x,r))\leq Cr^{\alpha},\;\;\;x\in A,0<r<\text{diam}(A)$
We say that $E$ is $\alpha$-AD regular if $E$ is $(\alpha,C)$-AD regular, for some constant $C\geq 1$.
The pointwise analog of AD-regular sets is defined in the natural way. That is, a point $x\in\mathbb{R}^{n}$ is $(\alpha,C)$-AD regular with respect to an oracle $A\subseteq\mathbb{N}$ if
(34) $\alpha r-C\log r\leq K_{r}^{A}(x)\leq\alpha r+C\log r$
for every $r\in\mathbb{N}$.
Let $E\subseteq\mathbb{R}^{n}$ be a compact set which is $\alpha$-AD regular. We say that an oracle $A\subseteq\mathbb{N}$ is an AD-regular Hausdorff oracle, or AD-regular oracle, for $E$ if
(35) $\alpha r-O(\log r)\leq K_{r}^{A}(x)\leq\alpha r+O(\log r)\;\forall r\in\mathbb{N}$
for $\mathcal{H}^{\alpha}$-a.e. $x\in E$. Note that we immediately have that any AD-regular oracle for $E$ is a Hausdorff oracle for $E$.
Lemma 42.
Let $E\subseteq\mathbb{R}^{n}$ be an $\alpha$-AD regular set. Then there is an AD-regular oracle $A$ for $E$. Moreover, for any oracle $B\subseteq\mathbb{N}$, the join $(A,B)$ of $A$ and $B$ is also an AD-regular oracle for $E$.
Proof.
Let $A\subseteq\mathbb{N}$ be an oracle relative to which $E$ is effectively compact, which encodes the real $\alpha$, and which encodes the set of reals
$\{\mathcal{H}^{\alpha}(E\cap Q)\mid Q\in\mathcal{Q}\}$,
where $\mathcal{Q}$ is the set of balls with rational centers and rational radii.
Since the restriction $\mu:=\mathcal{H}^{\alpha}|_{E}$ is a computable outer measure relative to $A$, and $0<\mu(E)<\infty$, by [27][Lemma 36], the set
$\{x\in E\mid(\exists^{\infty}r)\;K^{A}_{r}(x)<\alpha r-O(\log r)\}$
has $\mu$-measure zero.
Since $E$ is $(\alpha,C)$-AD regular, there is a constant $C^{\prime}>1$ such that, for every $r\in\mathbb{N}$ there is a cover of $E$ by rational balls of radius $2^{-r}$ of cardinality at most $C^{\prime}2^{\alpha r}$. Moreover, since $E$ is computably compact relative to $A$, this cover is computable relative to $A$. Therefore, for every $x\in E$ and every $r\in\mathbb{N}$,
$K^{A}_{r}(x)\leq\alpha r+O(\log r)$,
and the proof is complete. ∎
Theorem 43.
Suppose that $x,y\in\mathbb{R}^{2}$, $e=\frac{y-x}{|y-x|}$, $\sigma\in\mathbb{Q}\cap(0,1)$, $d_{x}>\sigma$, $\varepsilon>0$, $c\in\mathbb{N}$ and $A\subseteq\mathbb{N}$ satisfy the following.
  • (C1)
    $\dim^{A}(x)=d_{x}>1,$ $1\geq\dim^{A}(y)>\sigma$
  • (C2)
    $K^{A,x}_{r}(e)>\sigma r-O(\log r)$ for all sufficiently large $r$.
  • (C3)
    $K^{A,x}_{r}(y)\geq K^{A}_{r}(y)-O(\log r)$ for all sufficiently large $r$.
  • (C4)
    $K^{A}_{t,r}(e\mid y)>\sigma t-O(\log r)$ for all sufficiently large $r$ and $t\leq r$.
  • (C5)
    $d_{x}r-c\log r\leq K_{r}^{A}(x)\leq d_{x}r+c\log r$
Then
$\dim^{x,A}(|x-y|)\geq\sigma.$
Proof.
Note that condition (C5) immediately implies that $\operatorname{Dim}^{A}(x)-\dim^{A}(x)=0$. Therefore, by Theorem 39, for every $\varepsilon>0$,
(36) $\dim^{A,x}(|x-y|)\geq\sigma-\frac{10\varepsilon}{\sigma},$
and the conclusion follows. ∎
Theorem 44.
Any compact AD-regular $X\subseteq\mathbb{R}^{2}$ of dimension more than 1 is universal for pinned distances for the class of Borel sets.
Proof.
Assume $X\subseteq\mathbb{R}^{2}$ is compact and AD regular with dimension $d_{x}>1$ and regularity oracle $A_{X}$. Let $\mu$ be the normalized probability on $X$.
Let $Y\subseteq\mathbb{R}^{2}$ Borel of Hausdorff dimension at most $1$ be given, and define $Y_{i}$ to be a sequence of compact subsets of $Y$ satisfying $0<\mathcal{H}^{\dim_{H}(Y)-\frac{1}{i}}(Y_{i})<\infty$. We note the probability measures on $Y_{i}$ (the normalized $\mathcal{H}$ measure) by $\nu_{i}$. Let $A_{i}$ be an effective compactness oracle, encoding $\nu_{i}$, for $Y_{i}$. Define $A$ to be the join of $A_{X}$ and $A_{1},A_{2},...$ and observe that $A$ remains an AD-regularity oracle for $X$ and is an effective compactness oracle for all $Y_{i}$. Define
$X_{1}=\{x\in X:(\exists c)\,(\forall r)\text{ }d_{x}r-c\log r\leq K_{r}^{A}(x)\leq d_{x}r+c\log r\}$
Note that $\mu(X_{1})=1$.
For each $i$, let $\sigma_{i}=\dim_{H}(Y)-\frac{2}{i}$. Let $i\in\mathbb{N}$. Note that, if $\nu_{i}(\ell)\neq 0$, for any line $\ell\subseteq\mathbb{R}^{2}$, then $\dim_{H}(\Delta_{x}Y_{i})=\dim_{H}(Y_{i})$, for almost every $x\in X$. Without loss of generality, we assume that $\nu_{i}(\ell)=0$ for every line $\ell\subseteq\mathbb{R}^{2}$. By Lemma 2.1 of [21], we have that, for any $\varepsilon>0$, there is a $\delta>0$ such that $\nu_{i}(T)\leq\varepsilon$ for all $\delta$-tubes $T$.
We may therefore apply Corollary 2.18 of [22] to conclude that, for some $K_{i}$, $(\mu,\nu_{i})$ has $(\sigma_{i}^{\prime},K_{i},1-2^{-2i})$-thin tubes, for every $\sigma_{i}<\sigma^{\prime}_{i}<\dim_{H}(Y_{i})$. In particular, there is an $x\in X_{1}$ such that for every $i\in\mathbb{N}$, there is a point $y\in Y_{i}$ such that $x,y$ and $e:=\frac{y-x}{|x-y|}$ satisfy (C1)-(C3) and (C5). Using Lemma 34, we see that they also satisfy (C5). We may therefore apply Theorem 43 and conclude that
$\dim^{A,x}(|x-y|)\geq\dim_{H}(Y)-\frac{2}{i}$.
Since $Y_{i}$ is effectively compact relative to $A$, this implies that
(37) $\dim_{H}(\Delta_{x}Y_{i})\geq\dim_{H}(Y)-\frac{2}{i},$
and the conclusion follows. ∎

### v2

@@
 $\dim_{H}(\Delta_{x}Y)=\min\{\dim_{H}(Y),1\}$.
+2020 Mathematics Subject Classification:
+28A78, 28A80, 68Q30
 The first author was supported in part by NSF DMS-2037851 and NSF DMS-2246906.
@@
 $\Delta(E):=\{|x-y|\mid x,y\in E\}$.
-A fundamental problem in geometric measure theory is to relate the (Hausdorff) dimension of $E$ and the dimension of its distance set $\Delta(E)$.
-Falconer conjectured [7] that, if $E$ is analytic and $\dim_{H}(E)>n/2$, then $\Delta(E)$ has positive Lebesgue measure (and so $\dim_{H}(\Delta(E))=1$).
-This is still unproven in all dimensions $n\geq 2$.
-A generalization of this is the pinned distance problem. If $E\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, the pinned distance set of $E$ with respect to $x$ is the set
+A fundamental problem is to relate the sizes of $E$ and $\Delta(E)$; if $E$ is large, it should contain “many” different distances.
+For finite sets $E$, this is the Erdős distinct distances problem, which was essentially solved in the plane by Guth and Katz [10].
+In the continuum setting, Falconer conjectured [7] that, if $E$ is Borel and has Hausdorff dimension more than $n/2$, then $\Delta(E)$ has positive Lebesgue measure.
+Falconer’s conjecture is still open in all dimensions $n\geq 2$.
+For $E$ no larger than the $n/2$ threshold, the distance sets do not necessarily have positive measure, but one still expects there to be some lower bound on the Hausdorff dimension of $\Delta(E)$.
+Working in the plane, we improve upon the best known lower bounds of this type when .$1.5<\dim_{H}(E)<$ 1.
+In particular, we strengthen bounds on the size of “pinned” distance sets.
+If $E\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, the pinned distance set of $E$ with respect to $x$ is
 $\Delta_{x}(E):=\{|x-y|\mid y\in E\}$.
-The dimension version of Falconer’s conjecture for pinned distance sets is the following.
+Note that if $x\in E$, then $\Delta_{x}(E)\subseteq\Delta(E)$, so appropriate lower bounds on the size of pinned distance sets imply the same lower bounds for distance sets.
+The dimension version of Falconer’s conjecture for pinned distance sets is the following:
 Conjecture 1.
-Let $E\subseteq\mathbb{R}^{n}$ be an analytic set with $\dim_{H}(E)>n/2$. Then
+Let $E\subseteq\mathbb{R}^{n}$ be a Borel set with $\dim_{H}(E)>n/2$. Then
 $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))=1$.
-Substantial progress has been made on this question in the past few years [6, 9, 3, 4, 5].
-A full resolution of Conjecture 1 problem seems to be beyond current techniques, even in the $n=2$ case.
-However, recent work has proven strong lower bounds on the dimension of pinned distance sets in the plane [23, 24, 11, 27].
-If $E\subseteq\mathbb{R}^{2}$ is analytic, and $d:=\dim_{H}(E)>1$, the current records are
+Although a full resolution of Conjecture 1 seems to be beyond current techniques, substantial progress has been made on this question in the past few years [6, 9, 3, 4, 5].
+Moreover, recent work has proven strong lower bounds on the dimension of pinned distance sets in the plane [25, 26, 12, 29].
+If $E\subseteq\mathbb{R}^{2}$ is analytic (i.e.
+the continuous image of a Borel set), and $d:=\dim_{H}(E)>1$, the current records are
 […]
-Unfortunately, the techniques in both [12] and [8] rely crucially on the assumption that $\dim_{H}(E)>1$.
-In this paper, we prove lower bounds on the pinned distance sets of planar sets of small dimension.
-There is significantly less known in this setting.
-A fundamental issue had been the lack of progress on radial projections for low-dimensional sets.
-Nevertheless, Shmerkin and Wang gave the first strong progress on this problem [^1] A key ingredient of Shmerkin and Wang’s proof was a result on radial projections of low dimensional sets.
-This component was then improved in [22]..
-They proved [25] that, for any analytic set $E\subseteq\mathbb{R}^{2}$ with $d:=\dim_{H}(E)\leq 1$,
-(1) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{d}{2}+\frac{d^{2}}{2(2+\sqrt{d^{2}+4})}.$
-Moreover, they proved that if $E$ is weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$, then $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)=d$. In [5], Du, Ou, Ren and Zhang proved that
-(2) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{5}{3}d-1,$
-which improves on (1) when $d$ is sufficiently close to one.
-In this paper, we establish stronger bounds on the dimension of pinned distance sets, for certain ranges of $d$. Our main theorem is the following.
+However, the techniques in both [13] and [8] rely crucially on the assumption that $\dim_{H}(E)>1$.
+In this paper, we establish improved bounds on the dimension of the pinned distance sets in the absence of this assumption.
+Let $\dim_{P}(E)$ denote the packing dimension of $E$ (another notion of fractal dimension that is dual to Hausdorff dimension and always satisfies $\dim_{P}(E)\geq\dim_{H}(E)$).
+Our main theorem is the following:
 Theorem 2.
-Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set, $d:=\dim_{H}(E)\leq 1$ and $D:=\dim_{P}(E)$. Then,
+Suppose $E\subseteq\mathbb{R}^{2}$ is analytic, $d:=\dim_{H}(E)\leq 1,$ and $D:=\dim_{P}(E)$. Then,
 $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{\alpha D-d(D+\alpha-d)}{(d+1)(\alpha D-d^{2})-d^{2}(\alpha+D-2d)}\right)$,
 where $\alpha=\min\{1+d,D\}$.
-Note that the bound improves as $E$ becomes more regular, i.e., as the packing dimension of $E$ approaches its Hausdorff dimension.
-An important special case (when $D=2$) is captured by the following corollary.
+Note that the bound improves as $E$ becomes more regular, i.e., as the packing dimension of $E$ approaches its Hausdorff dimension. Hence, setting $D=2$ gives the following corollary.
 Corollary 3.
-Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set and $d:=\dim_{H}(E)\leq 1$. Then,
+Suppose $E\subseteq\mathbb{R}^{2}$ is analytic and $d:=\dim_{H}(E)\leq 1$. Then,
 $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}(E))\geq d\left(1-\frac{2-d}{2(1+2d-d^{2})}\right)$.
-The proof of this theorem uses effective methods, and follows the general framework of the authors previous paper [8].
-The assumption that $\dim_{H}(E)<1$ necessitates important changes to previous work.
+Until recently, less progress had been made on the study of pinned distance sets in this setting than above the $n/2$ threshold.
+A fundamental issue had been the need for improved bounds on the size of radial projections for small sets.
+Shmerkin and Wang gave the first strong progress on this problem, proving [27] that, for any analytic set $E\subseteq\mathbb{R}^{2}$ with $d:=\dim_{H}(E)\leq 1$,
+(1) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{d}{2}+\frac{d^{2}}{2(2+\sqrt{d^{2}+4})}.$
+Moreover, they proved that if $E$ is weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$, then $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)=d$.
+A key ingredient of Shmerkin and Wang’s proof was a result on radial projections of low-dimensional sets which was then improved in [24].
+In [5], Du, Ou, Ren and Zhang proved that
+(2) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{5}{3}d-1,$
+which exceeds (1) when $d$ is sufficiently close to one.
+Corollary 3 improves on both of these bounds provided that $d$ is greater than about .477.477.477.477.
+Furthermore, the lower bound is $3/4$ when $d=1$, which agrees with the lower bound in [8] and eliminates the discontinuity that had existed at this threshold value.
+The proof of our main theorem proceeds via “effective” methods.
+In particular, we study the algorithmic information content of points as well as the distances and directions between them.
+Appropriate bounds on the complexity of distances $|x-y|$ imply bounds on the size of the pinned distance set $\Delta_{x}(E)$ via the point-to-set principle of Lutz and Lutz [15].
+The proof broadly follows a similar framework to the authors’ previous paper [8].
+However, the assumption that $\dim_{H}(E)\leq 1$ necessitates new techniques relative to previous work.
 One of the most crucial differences is the use of the recent breakthrough by Orponen, Shmerkin and Wang [24] on radial projections for low-dimensional sets.
 In the language of effective dimension, this allows us to find $x,y\in E$ whose direction $e:=\frac{y-x}{|x-y|}$ is sufficiently complex.
-Additionally, a number of the algorithmic tools of [8] need to be adapted to this setting.
+Making use of this theorem requires a new reduction, i.e.
+a proof that the main theorem follows from its effective analog.
+Additionally, versions of many of the algorithmic tools of [8] need to be established in this setting, and a careful usage of these tools requires leveraging certain bounds only relevant when the information content of objects is low, as in the case of small sets.
+1.1. Universal sets for pinned distances
 The next main theorems of this paper deal with the concept of universal sets.
@@
 (3) $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))=\min\{1,\dim_{H}(Y)\}.$
+Note that the right hand side is as large as the pinned distance set can possibly be in the sense of Hausdorff dimension.
 We can strengthen this notion by requiring the existence of specific points in $X$ achieving the desired bound.
@@
 Unpacking the definitions, this theorem shows that there is a fixed set $X\subseteq\mathbb{R}^{2}$ such that, for every $Y$ with $\dim_{H}(Y)<\dim_{H}(X)$, $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))$ is maximal.
-Moreover, combined with a previous theorem of the authors [8], we have the following corollary.
+This can be thought of as strengthening the result in [27] that if $Y$ is weakly regular, then it has pinned distance sets of the maximum possible size.
+We are able to obtain the same conclusion using only regularity in the pin set $X$.
+Moreover, combined with a previous theorem of the authors (Theorem 7 in [8], which handles the case that $Y$ is large), we have the following corollary.
 Corollary 5.
@@
 Let $X\subseteq\mathbb{R}^{2}$ be a compact, Ahlfors-David regular set such that $\dim_{H}(X)>1$. Then $X$ is universal for pinned distances for the class of Borel sets $Y\subseteq\mathbb{R}^{2}$.
-In particular, if we take any self-similar four-corner Cantor set $C$ of dimension strictly greater than $1$, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in C$ such that
+For instance, if we take any self-similar four-corner Cantor set $C$ of dimension strictly greater than $1$, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in C$ such that
 $\dim_{H}(\Delta_{x}Y)=\min\{\dim_{H}(Y),1\}.$
+Considering again the “unpinned” distance set problem gives another interpretation of Theorem 6. Sets such as $C$ have the property that for any Borel $Y$, one can choose $x\in C$ such that
+$\dim_{H}(\Delta(Y\cup\{x\}))\geq\min\{\dim_{H}(Y),1\}$
+In particular, (the dimension version of) Falconer’s conjecture is true in the plane for any Borel $Y$, provided that we are allowed to add one point drawn from a small fixed set to $Y$.
 2. Preliminaries
@@
 Case $S=1$: Suppose the green portion of our red-green-blue sequence starts at precision $r_{1}$ and ends at precision $r_{1}+s$, for some $s\geq t$.
-Let $r_{2}\geq r_{1}+s$ be the minimal precision such that $[r_{2},r]$ is the union of yellow intervals [^2] No such precision necessarily exists.
-In this case, set $r_{2}=r$..
+Let $r_{2}\geq r_{1}+s$ be the minimal precision such that $[r_{2},r]$ is the union of yellow intervals; if no such precision exists, then set $r_{2}=r$.
 We first assume that $B>\frac{\sigma}{\alpha}(r-t)$, where $B$ is the total length of the non-teal intervals in $\mathcal{P}$. Then, by Lemma 18, we have
@@
 We note that $e^{\bot}$ and $e$ are Lipschitz-bi-computable, and so $K_{r}(e)=K_{r}(e^{\bot})+O(\log r)$.
-By a pointwise analog of Marstrand’s projection theorem, [^3] Essentially proven in [17], but also a result of the projection bound in section 3 arising from a one interval partition
+By a pointwise analog of Marstrand’s projection theorem (which was essentially proven in [19], but is also a result of the projection bound in section 3 arising from a one interval partition), we have
 $K^{A}_{t}(x\mid p_{e^{\bot}}x,e^{\bot})<K^{A}_{t}(x)-\sigma t+O(\log t),$
@@
 This is the content of the first subsection.
-In the second subsection, we assume additional regularity in $X$, namely that it is AD regular [^4] Full Ahlfors-David regularity isn’t necessary, it is possible to weaken the assumption slightly.
-However, we will require AD regularity to simplify the exposition..
-In this case, we are guaranteed the existence of points that are extremely regular, with their complexity functions only deviating from $\dim_{H}(x)r$ by a logarithmic term.
+In the second subsection, we assume additional regularity in $X$, namely that it is AD regular.
+Full Ahlfors-David regularity isn’t necessary; it is possible to weaken the assumption slightly.
+However, we will require AD regularity to simplify the exposition.
+In the AD regular case, we are guaranteed the existence of points that are extremely regular, with their complexity functions only deviating from $\dim_{H}(x)r$ by a logarithmic term.
 This enables us to show an optimal effective theorem.